### v3

# Jaime Angulo Pava  and  Alexander Muñoz

Abstract
The aim of this work is to study the Airy and Schrödinger operators on looping-edge graphs, a graph consisting of a circle and a finite number $N$ of infinite half-lines attached to a common vertex. Characterizations of extensions generating unitary and contractile dynamics are provided for the Airy operator in terms of self-orthogonal subspaces and linear operators acting on indefinite inner product spaces (Krein spaces) associated to the boundary values at the vertex. Employing the same abstract techniques used in the former case, we characterize the self-adjoint extensions of the Schrödinger operator on $\mathcal{T}$-shaped graphs, i.e., to graphs that consists of an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with $\nu=0$. The theory established in this manuscript can be considered as a first step towards a better understanding of the dynamics for the remarkable dispersive models NLS and KdV on metric graphs.
Mathematics Subject Classification (2020). Primary 35Q51, 35Q55, 81Q35, 35R02; Secondary 47E05.
Key words. Schrödinger equation, Airy equation, quantum graphs, Krein spaces, extension theory of symmetric operators.
1. Introduction
Nonlinear evolution models on metric graphs arise as models in wave propagation, for instance, in quasi-one-dimensional systems (e.g., meso- or nanoscale) that resemble a thin neighborhood of a graph. These models provide a convenient means to study various physical effects in the real world, both from theoretical and experimental perspectives. The flexibility in setting the geometry of the graph configuration allows for the creation of different dynamics. To mention a few examples of bifurcated systems: Josephson junction structures, networks of planar optical waveguides and fibers, branched structures associated with DNA, blood pressure waves in large arteries, nerve impulses in complex arrays of neurons, conducting polymers, and Bose-Einstein condensation (see [17, 19, 20, 24, 25, 27, 30, 35, 37, 40, 45] and reference therein).
Recently, nonlinear models on graphs, such as the nonlinear Schrödinger equation, the sine-Gordon model, and the Korteweg-de Vries model, have been studied extensively (see [2, 11, 12, 8, 9, 13, 14, 15, 3, 30, 40] and references therein). From a mathematical viewpoint, an evolution model on a graph is equivalent to a system of PDEs defined on the edges (intervals), where the coupling is determined exclusively by the boundary conditions at the vertices (known as the ”topology of the graph”), which governs the dynamics on the network via groups. This area of study has attracted significant attention, particularly in the context of soliton transport. Solitons and other nonlinear waves in branched systems provide valuable insights into the dynamics of these models.
Addressing these issues is challenging because both the equations of motion and the graph topology can be complex. Additionally, a critical aspect that complicates the analysis is the presence of one or more vertices where a soliton profile, arriving at the vertex along one of the edges, undergoes complicated behaviors such as reflection and the emergence of radiation. This makes it difficult to observe how energy travels across the network. As a result, the study of soliton propagation through networks presents significant challenges. The mechanisms for the existence and stability (or instability) of soliton profiles remain unclear for many types of graphs and models.
We recall that a metric graph $\mathcal{G}$ is a structure represented by a finite number of vertices $V=\{\nu_{i}\}$ and a set of adjacent edges at the vertices $E=\{e_{j}\}$ (for further details, see [17]). Each edge $e_{j}$ can be identified with a finite or infinite interval of the real line, $I_{e}$. Thus, the edges of $\mathcal{G}$ are not merely abstract relations between vertices but can be viewed as physical ”wires” or ”networks” connecting them. The notation $e\in E$ will be used to indicate that $e$ is an edge of $\mathcal{G}$. This identification introduces the coordinate $x_{e}$ along the edge $e$.
An example of a metric graph, which will be the focus of our study here, is a looping-edge graph. This is a graph composed of a circle and a finite number $N$ of infinite half-lines attached to a common vertex $\nu=L$. If the circle is identified with the interval $[-L,L]$ and the half-lines with $[L,\infty)$, we obtain a particular metric graph structure that we denote with $\mathcal{G}$, represented by $E=\{[-L,L],[L,\infty),\dots,[L,\infty)\}$ and $V=\{L\}$ (see Figure 1). We use the subscripts 00 and $j$ to refer to the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$, $j=1,\dots,N$, respectively. The special case $N=1$ is known as a tadpole graph (see also Figure 1).
$-L$
$L$
$-L$
$L$
Tadpole graph
Looping-edge graph
Figure 1. Tadpole and looping-edge-type graphs
A wave function U defined on $\mathcal{G}$ will be understood as an $(N+1)$-tuple of functions $\textbf{U}=(\phi,(\psi_{j})_{j=1}^{N})=(\phi,\psi_{1},\dots,\psi_{N})$, where $\phi$ is defined on $e_{0}=[-L,L]$ and $\psi_{j}$ is defined on $e_{j}=[L,\infty)$ for $j=1,\dots,N$.
Consider the case of nonlinear Schrödinger-type equations (NLS)
$i\textbf{U}_{t}\ +\Delta\textbf{U}+(p+1)|\textbf{U}|^{2p}\textbf{U}=\textbf{0},\quad p>0,$ (1.1)
where the action of the Laplacian operator $\Delta$ on a general graph $\mathbb{G}$ is given by
$-\Delta:(u_{e})_{e\in E}\to(-u^{\prime\prime}_{e})_{e\in E}.$ (1.2)
The NLS in (1.1) has been extensively studied in recent literature for various types of metric graphs $\mathbb{G}$ and with specific domains for the Laplacian, which make it a self-adjoint operator on $L^{2}(\mathbb{G})$ (see the review manuscript [34]). For instance, research has focused on star graphs ([2, 1, 11, 12, 32, 40] and references therein), looping-edge graphs ([6, 7, 21, 42, 41] and references therein), as well as flower graphs, dumbbell graphs, double-bridge graphs, and periodic ring graphs ([21, 18, 33, 31, 36, 41, 43] and references therein).
One of the objectives of this work is to obtain a characterization of all self-adjoint extensions of the symmetric operator $\mathcal{H}_{0}\equiv-\Delta$ on the metric graph $\mathcal{G}$ with domain $D(\mathcal{H}_{0})$ defined by
$D(\mathcal{H}_{0})=C_{0}^{\infty}(-L,L)\oplus\bigoplus_{j=1}^{N}C_{0}^{\infty}(L,+\infty).$ (1.3)
Since the deficiency indices of $(\mathcal{H}_{0},D(\mathcal{H}_{0}))$ are equal to $2+N$, the extension theory of Krein and von Neumann provides a $(2+N)^{2}$-parameter family of self-adjoint extensions $(-\Delta_{ext},D(-\Delta_{ext}))$ of $\mathcal{H}_{0}$ on $L^{2}(\mathcal{G})$, where the action of $-\Delta_{ext}$ is given by $-\Delta_{ext}=-\Delta$. Each member of this family defines a unitary dynamic for the linear evolution problem
$\left\{\begin{array}[]{ll}i\textbf{U}_{t}=-\Delta_{ext}\textbf{U}\\ \textbf{U}(0)=\textbf{u}_{0}\in D(-\Delta_{ext}).\end{array}\right.$ (1.4)
We point out that our approach, based on Krein spaces [44], demonstrates that every extension operator $(-\Delta_{ext},D(-\Delta_{ext}))$ can be characterized by the boundary conditions at the vertices $\nu=\pm L$, treating the point $-L$ as a terminal vertex through a relatively simple matrix process (see Theorems 5.2-5.3 below). At this level, our approach can be seen as a complement to the works of Exner et al. [27, 28, 29], which, to our knowledge, has not formally appeared in the literature on $\mathcal{G}$. Furthermore, in general, the extensions $(-\Delta_{ext},-D(\Delta_{ext}))$ do not necessarily satisfy continuity conditions at the boundary of $[-L,L]$. (see, for instance, Example 8 associated to $\delta^{\prime}$- type interactions on looping-edge graphs and Example 10 below by considering $-L$ as a terminal vertex). Moreover, our approach to obtaining these extensions will be based on those associated with the Airy operator (1.7) which is the main focus of our study. Both extensions theories to be established can be seen as a first step for future studies on the existence and stability of soliton profiles for both NLS-(1.1) and KdV-(1.8) models, for instance, on a looping-edge graph or on a $\mathcal{T}$-shaped graph, i.e., a graph that consists of an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with $\nu=0$ (see Figure 4). In the case of looping-edge graphs, the extensions must ensure that the wave functions are continuous at least on the loop (i.e., $\phi(L)=\phi(-L)$). For example, in the study of electron motion in thin metallic ”wires,” it is required that the wave function remains continuous at the junction $\nu=L$. These boundary conditions have a clear physical interpretation, representing the conservation of probability flow at the junction (see [28]).
This physical requirement can be formally expressed through the following family of self-adjoint operator extensions $(-\Delta,D_{Z,N})_{Z\in\mathbb{R}}$ for $(\mathcal{H}_{0},D(\mathcal{H}_{0}))$ with
$\begin{array}[]{ll}D_{Z,N}=\{\mathbf{U}\in H^{2}(\mathcal{G}):&\phi(L)=\phi(-L)=\psi_{1}(L)=\cdot\cdot\cdot=\psi_{N}(L)\;\text{and}\\ &\phi^{\prime}(L)-\phi^{\prime}(-L)=\sum_{j=1}^{N}\psi_{j}^{\prime}(L+)+Z\psi_{1}(L)\},\end{array}$ (1.5)
where for any $n\geqq 0$,
$H^{n}(\mathcal{G})=H^{n}(-L,L)\oplus\bigoplus_{j=1}^{N}H^{n}(L,+\infty).$
The boundary conditions in (1.5) are called $\delta$-type if $Z\neq 0$, and Neumann-Kirchhoff type if $Z=0$. The parameter $Z$ is a coupling constant between the loop and the several half-lines (see Example 7 and Remark 1). The choice of coupling at the vertex $\nu=L$ corresponds to a feasible quantum-wire experiment (see [27], [29], and references therein). It should be noted that, based on the domain $D_{Z,N}$, several recent studies have been conducted, particularly on the existence and stability of standing-wave solutions for the model in (1.1) (see [3, 4, 5, 6, 7, 16, 21, 22, 34, 42, 41], and references therein).
Our second focus in this manuscript is associated to obtain a characterization of all boundary conditions under which the Airy-type evolution equation
$\partial_{t}u_{\textbf{e}}(x,t)=\alpha_{\textbf{e}}\partial_{x}^{3}u_{\textbf{e}}(x,t)+\beta_{\textbf{e}}\partial_{x}u_{\textbf{e}}(x,t),\quad x\neq 0,t\in\mathbb{R},\textbf{e}\in E,$ (1.6)
with real coefficients given by the sequences $(\alpha_{\textbf{e}})_{e\in E}$, $(\beta_{\textbf{e}})_{e\in E}$, generates either a contraction semigroup or a unitary group on a looping edge graph (see Theorems 4.2, 4.3 and 4.7). Our approach is based on the abstract results in [44] applied to the case of the Airy-operator
$A_{0}:\{u_{e}\}_{e\in E}\mapsto\{\alpha_{e}\partial_{x}^{3}u_{e}+\beta_{e}\partial_{x}u_{e}\}_{e\in E},$ (1.7)
when we consider the metric graph $\mathcal{G}$ represented by $E$ defined above and treating the point $-L$ as a terminal vertex (as in the case of a $\mathcal{T}$-graph). Therefore, notions of self-orthogonal subspaces and operators based on indefinite inner product spaces (Krein spaces) will be necessary. Thus, we get in particular a description of all (skew-)self-adjoint realizations of $A_{0}$, - i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$. For instance, in Example 1 below we get $\delta$-interactions type boundary condictions on a looping-edge graph in a similar way as done in the case of metric star-shaped graphs (see, for instance, Theorem 3.6 in [8]) under conditions on $\alpha_{\textbf{e}}$ and $\beta_{\textbf{e}}$. Also, we establish conditions for obtaining extensions of $A_{0}$ generating contractile dynamics.
We recall that the evolution linear model in (1.6) is associated to the following vectorial Korteweg-de Vries (KdV) equation
$\partial_{t}u_{\textbf{e}}(x,t)=\alpha_{\textbf{e}}\partial_{x}^{3}u_{\textbf{e}}(x,t)+\beta_{\textbf{e}}\partial_{x}u_{\textbf{e}}(x,t)+2u_{\textbf{e}}(x,t)\partial_{x}u_{\textbf{e}}(x,t),$ (1.8)
on the metric graph $\mathcal{G}$. Recently, Mugnolo, Noja, and Seifert [38] characterized all boundary conditions under which the Airy equation in (1.6) generates either a semigroup or a group on a metric graph that has a structure represented by the set $E\equiv E_{-}\cup E_{+}$, where ${E}_{+}$ and ${E}_{-}$ are finite or countable collections of semi-infinite edges $e$ parametrized by $(-\infty,0)$ and $(0,+\infty)$, respectively. The half-lines are connected at a unique vertex $\nu=0$. In this case the graph is sometimes called a star-shaped metric graph. Several arguments used in [44] and [38] in star-shaped graphs are extended in this work for looping-edge graphs. We recall that when the star-shaped metric graph is balanced ($|{E}_{+}|=|{E}_{-}|$), the extension theory implies that there is a $9|{E}_{+}|^{2}$-parameter family of skew-self-adjoint operators for $A_{0}$. Now, on looping-edge graph when the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we get that there must be $9(k+1)^{2}$ family of unitary groups, each one having its own representation.
As far as we know, the study of the KdV equation on metric graphs remains relatively underdeveloped. Cavalcante [23] studied the local well-posedness of the Cauchy problem associated with (1.8) in Sobolev spaces $H^{s}(\mathcal{Y})$ with low regularity on a $\mathcal{Y}$-junction graph, $\mathcal{Y}=(-\infty,0)\cup(0,+\infty)\cup(0,+\infty)$. Subsequently, Angulo and Cavalcante [8] established a criterion for the linear instability of stationary solutions for the KdV on arbitrary star-shaped metric graphs. This criterion was applied to tail and bump profiles on balanced graphs with specific boundary conditions at the vertex $\nu=0$, ensuring they remain linearly unstable (see Theorem 6.2 in [8]). More recently, Mugnolo, Noja, and Seifert [39] investigated the existence of solitary waves on a star-shaped metric graph in terms of the coefficients of the KdV equation $(\alpha_{\mathbf{e}})_{e\in E}$, $(\beta_{\mathbf{e}})_{e\in E}$ on each edge, with coupling conditions at the vertex $\nu=0$, which may or may not involve continuity, and examined the speeds of the traveling waves. Lastly, Angulo and Cavalcante [10] developed a local well-posedness theory for the KdV on a balanced star graph, considering boundary conditions at the vertex $\nu=0$ of $\delta$- interaction type. They also extended the linear instability result from [8] to establish nonlinear instability.
Results of the dynamics of the KdV model on looping edge graphs or on $\mathcal{T}$-shaped graphs, such as local well-posedness of the Cauchy problem, existence and stability of solitary waves, etc., are not known in the literature to the best of the authors knowledge. We believe that the theory established in this manuscript can be considered as a first step towards a better understanding of this notable dispersive model on metric graphs. Organization of the article.
  1. 2.
    Preliminaries
    1. 2.1.
      Notation
    2. 2.2.
      Stone and Lummer-Phillips Theorems
    3. 2.3.
      Operators on Krein Spaces
  2. 3.
    The free Airy operator and deficiency indices
  3. 4.
    The Airy operator on a looping-edge graph
    1. 4.1.
      Boundary system for even amount of half-lines
      1. 4.1.1.
        Extensions generating unitary dynamics
      2. 4.1.2.
        Examples: $\delta$-type interactions on a looping-edge graph
    2. 4.2.
      Boundary system for arbitrary amount of half-lines
      1. 4.2.1
        Extensions generating contractile dynamics
      2. 4.2.2
        Examples: $\delta$-type interactions on a tadpole graph
    3. 4.3.
      Boundary system separating derivatives
      1. 4.3.1
        Extensions generating contractile dynamics via splitting derivatives
      2. 4.3.2
        Examples
  4. 5.
    The Schrödinger operator on looping-edge and $\mathcal{T}$-shaped metric graphs.
    1. 5.1.
      Deficiency indices for the free Schrödinger operator
    2. 5.2.
      Boundary system
    3. 5.3.
      Extensions generating unitary dynamics
    4. 5.4.
      Extensions generating unitary dynamics parameterized by linear subspaces
    5. 5.5.
      Extensions to $\mathcal{T}$-shaped metric graphs
2. Preliminaries
2.1. Notation
Let us introduce some notation related to the boundary values of a function posed in $\mathcal{G}$. For $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$, we denote the vector of values of U and its derivatives at the vertex with
$\vec{\textbf{U}}:=\begin{pmatrix}\phi(-L)\\ \phi(L)\\ \psi_{1}(L)\\ \vdots\\ \psi_{N}(L)\end{pmatrix},\ \vec{\textbf{U}}^{\prime}:=\begin{pmatrix}\phi^{\prime}(-L)\\ \phi^{\prime}(L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mbox{ and }\vec{\textbf{U}}^{\prime\prime}:=\begin{pmatrix}\phi^{\prime\prime}(-L)\\ \phi^{\prime\prime}(L)\\ \psi_{1}^{\prime\prime}(L)\\ \vdots\\ \psi_{N}^{\prime\prime}(L)\end{pmatrix}.$
For the components of U, we denote the derivatives values at the vertex with
$\partial\phi(-L):=\begin{pmatrix}\phi(-L)\\ \phi^{\prime}(-L)\\ \phi^{\prime\prime}(-L)\end{pmatrix},\ \partial\phi(L):=\begin{pmatrix}\phi(L)\\ \phi^{\prime}(L)\\ \phi^{\prime\prime}(L)\end{pmatrix}\mbox{ and }\partial\psi_{j}(L)=\begin{pmatrix}\psi_{j}(L)\\ \psi_{j}^{\prime}(L)\\ \psi_{j}^{\prime\prime}(L)\end{pmatrix}.$
We adopt concatenation of vectors in vertical stack. For instance,
$\begin{pmatrix}\partial\phi(-L)\\ \partial\phi(L)\end{pmatrix}=\begin{pmatrix}\phi(-L)\\ \phi^{\prime}(-L)\\ \phi^{\prime\prime}(-L)\\ \phi(L)\\ \phi^{\prime}(L)\\ \phi^{\prime\prime}(L)\end{pmatrix}.$
For convenience we denote for $\psi=(\psi_{1},\dots,\psi_{N})$
$\partial\psi(L):=\begin{pmatrix}\partial\psi_{1}(L)\\ \partial\psi_{2}(L)\\ \vdots\\ \partial\psi_{N}(L)\end{pmatrix},\ \partial\psi_{even}(L):=\begin{pmatrix}\partial\psi_{2}(L)\\ \partial\psi_{4}(L)\\ \vdots\\ \partial\psi_{2k}(L)\end{pmatrix}\mbox{ and }\partial\psi_{odd}(L):=\begin{pmatrix}\partial\psi_{1}(L)\\ \partial\psi_{3}(L)\\ \vdots\\ \partial\psi_{2k-1}(L)\end{pmatrix}.$ (2.1)
2.2. Stone and Lumer-Phillips Theorems
In this subsections we recall some well-known features of group theory.
Let $X$ be a Banack space and $X^{\prime}$ its dual space. Consider $\mathcal{J}(x):=\{x^{\prime}\in X^{\prime}\mid\langle x,x^{\prime}\rangle=\|x\|^{2}=\|x^{\prime}\|^{2}\}$. Note $\mathcal{J}(x)$ is non-empty by Hahn-Banach Theorem. Moreover, if $X$ is a Hilbert space then $\mathcal{J}(x)$ consists of a single element.
Recall that a linear operator $(A,D(A))$ on a Banach space $X$ is called dissipative if $\|(\lambda-A)x\|\geqslant\lambda\|x\|$.
Proposition 2.1.
An operator $(A,D(A))$ is dissipative if and only if for every $x\in D(A)$ there exists $j(x)\in\mathcal{J}(x)$ such that
$\mbox{Re}\langle Ax,j(x)\rangle\leqslant 0.$ (2.2)
If $A$ is the generator of a strongly continuous contraction semigroup, then (2.2) holds for all $x\in D(A)$ and arbitrary $x^{\prime}\in\mathcal{J}(x)$.
Proof.
See Proposition 3.23 in [26]
Note in the case $X$ is a Hilbert space, condition (2.2) reduces to $\mbox{Re}(Ax\mid x)\leqslant 0$.
Theorem 2.2.
(Lummer-Phillips) Let $(A,D(A))$ be a densely defined operator on a Banach space $X$. If both, $A$ and $A^{*}$ are dissipative, then the closure $\overline{A}$ of $A$ generates a contraction semigroup on $X$.
Proof.
See Corollary 3.17 in [26]. ∎
We also recall the Stone’s theorem for unitary groups.
Theorem 2.3 (Stone).
Let $(A,D(A))$ be a densely defined operator on a Hilbert space $H$. Then $A$ generates a unitary group $(T(t))_{t\in\mathbb{R}}$ on $H$ if and only if $A$ is skew-self-adjoint, i.e., $A^{*}=-A$ and $D(A^{*})=D(A)$.
2.3. Operators on Krein Spaces
Following [44] we introduce the abstract notions of self-orthogonal subspaces and operators based on indefinite inner product spaces.
Definition 2.1.
Let $Y$ be a linear space, let $X\subset Y$ be a subspace and $w$ a sesquilinear form defined on $(Y\times Y)^{2}$. The space $X$ is called $w$-self-orthogonal if
$X=X^{\perp_{w}}\equiv\{(x,y)\in Y\times Y\mid w((x,y),(\tilde{x},\tilde{y}))=0\mbox{ for all }(\tilde{x},\tilde{y})\in X\}.$ (2.3)
Definition 2.2.
A Krein space $(\mathcal{K},(\cdot\mid\cdot)_{\mathcal{K}})$ is a complex linear space $\mathcal{K}$ on which a indefinite inner product $(\cdot\mid\cdot)_{\mathcal{K}}$ is defined, namely, the function $(x,y)\in\mathcal{K}\times\mathcal{K}\to(x\mid y)_{\mathcal{K}}\in\mathbb{C}$ satisfies the following axioms:
  1. i)
    Linearity in the first argument:
    $(\alpha x_{1}+\beta x_{2}\mid y)_{\mathcal{K}}=\alpha(x_{1}\mid y)_{\mathcal{K}}+\beta(x_{2}\mid y)_{\mathcal{K}}$
    for all $x_{1},x_{2},y\in\mathcal{K}$ and all complex numbers $\alpha,\beta$;
  2. ii)
    antisymmetry:
    $(x\mid y)_{\mathcal{K}}=\overline{(y\mid x)_{\mathcal{K}}}$
    for all $x,y\in\mathcal{K}$;
  3. iii)
    nondegeneracy: if $(x\mid y)_{\mathcal{K}}=0$ for all $y\in\mathcal{K}$, then $x=0$.
Thus, the function $(\cdot\mid\cdot)_{\mathcal{K}}$ satisfies all the properties of a standard inner product with the possible exception that $(x\mid x)_{\mathcal{K}}$ may be nonpositive for $x\neq 0$.
We recall that if $dim(\mathcal{K})=n$ ($\mathcal{K}\cong\mathbb{C}^{n}$) every indefinite inner product $(\cdot\mid\cdot)_{\mathcal{K}}$ on $\mathcal{K}$ is determined by an $n\times n$ invertible and hermitian matrix $H$ such that $(x\mid y)_{\mathcal{K}}=\langle Hx,y\rangle_{\mathbb{C}^{n}}$ holds for all $x,y\in\mathcal{K}$.
Definition 2.3.
Let $\mathcal{K}_{1},\mathcal{K}_{2}$ be Krein spaces and a linear operator $\mathcal{L}:\mathcal{K}_{1}\to\mathcal{K}_{2}$.
  1. (1)
    $\mathcal{L}$ is called a $(\mathcal{K}_{1},\mathcal{K}_{2})$-contraction if
    $\left(\mathcal{L}x\mid\mathcal{L}x\right)_{\mathcal{K}_{2}}\leqslant\left(x\mid x\right)_{\mathcal{K}_{1}}\mbox{ for all }x\in D(\mathcal{L}).$ (2.4)
  2. (2)
    If $\mathcal{L}$ is densely defined, the $(\mathcal{K}_{2},\mathcal{K}_{1})$-adjoint of $\mathcal{L}$ is defined by
    $D(\mathcal{L}^{\#}):=\{y\in\mathcal{K}_{2}\mid\exists z\in\mathcal{K}_{1}:\left(\mathcal{L}x\mid y\right)_{\mathcal{K}_{2}}=\left(x\mid z\right)_{\mathcal{K}_{1}}\mbox{ for all }x\in D(\mathcal{L})\}\mbox{ and }\mathcal{L}^{\#}y=z.$ (2.5)
  3. (3)
    $\mathcal{L}$ is called $(\mathcal{K}_{1},\mathcal{K}_{2})$-unitary if $D(\mathcal{L})$ and $R(\mathcal{L})$ are dense, $\mathcal{L}$ is injective and #$\mathcal{L}^{\#}=\mathcal{L}^{-1}$.
3. The free Airy operator and deficiency indices
In this section we fix notation and specify the functional spaces in which the Airy operator will be studied. We also provide an analysis on how the deficiency indices for the free Airy operator behave.
We consider the Airy operator
$A_{0}:\{u_{e}\}_{e\in E}\mapsto\{\alpha_{e}\partial_{x}^{3}u_{e}+\beta_{e}\partial_{x}u_{e}\}_{e\in E},$ (3.1)
where $\alpha_{j}$ and $\beta_{j}$ are real numbers for any $j$. For $\mathcal{G}$ determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering $-L$ as a terminal vertex, we understand the action of $A_{0}$ over the edges in a component-wise manner, that is, for $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$ posed on $\mathcal{G}$ we denote
$A_{0}\textbf{U}=\left(\alpha_{0}\partial_{x}^{3}\phi+\beta_{0}\partial_{x}\phi,\ \alpha_{1}\partial_{x}^{3}\psi_{1}+\beta_{1}\partial_{x}\psi_{1},\ \dots\ ,\ \alpha_{N}\partial_{x}^{3}\psi_{N}+\beta_{N}\partial_{x}\psi_{N}\right).$ (3.2)
We consider $A_{0}$ as a densely defined operator in
$L^{2}(\mathcal{G}):=L^{2}(-L,L)\oplus\bigoplus_{j=1}^{N}L^{2}(L,\infty)$ (3.3)
with domain
$D(A_{0})=C_{0}^{\infty}(-L,L)\oplus\bigoplus_{j=1}^{N}C_{0}^{\infty}(L,\infty).$ (3.4)
For $\textbf{U}=(\phi,\psi_{1}\dots,\psi_{N})$ and $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$, we denote the inner product on $L^{2}(\mathcal{G})$ with
[…]
We consider analogously the Sobolev spaces
$H^{s}(\mathcal{G}):=H^{s}(-L,L)\oplus\bigoplus_{j=1}^{N}H^{s}(L,\infty).$
Note in particular $D(A_{0}^{*})=H^{3}(\mathcal{G})$, where $A_{0}^{*}=-A_{0}$ is the adjoint of the minimal operator defined as the closure of $A_{0}$ with domain $D(A_{0})$ in (3.4).
For the densely defined symmetric operator $iA_{0}$ we consider the deficiency spaces
$\mathcal{D}_{\mp}(iA_{0}^{*}):=\mbox{Ker}(iA_{0}^{*}\pm iI),$ (3.6)
with deficiency indices $d_{\mp}=\mbox{dim}\mathcal{D}_{\mp}(iA_{0}^{*})$.
From the definition of $A_{0}$ in (3.1) we have that solutions of $-i\alpha_{e}\partial_{x}^{3}u_{e}-i\beta_{e}\partial_{x}u_{e}\pm iu_{e}=0$ without any boundary condition are obtained as linear combinations of complex exponentials of the form
$u_{e}=c_{1}e^{r_{1}x_{e}}+c_{2}e^{r_{2}x_{e}}+c_{3}e^{r_{3}x_{e}}$
where $r_{1},r_{2}$ and $r_{3}$ are the roots of $p(x)=\alpha_{e}x^{3}+\beta_{e}x\mp 1$. We refer the reader to the proof of Lemma 2.1 in [38] for an explicit description.
We note that $e^{rx}$ belong to $L^{2}(-L,L)$ independent on $r$, which means $d_{-_{0}}=d_{+_{0}}=3$ along $e_{0}$. On the other hand, on a half-line $[L,\infty)$ we note $e^{rx}$ is square integrable only if the real part of $r$ is negative, which depends on the signs of $\alpha_{e}$ and $\beta_{e}$. We summarize the behaviour (on a single edge $[L,\infty)$) in the following table.
Signs of $\alpha_{e}$ and $\beta_{e}$ $d_{-_{e}}$ $d_{+_{e}}$
$\alpha_{e}>0,\beta_{e}\geqslant 0$ 2 1
$\alpha_{e}>0,\beta_{e}<0$ 2 1
$\alpha_{e}<0,\beta_{e}\geqslant 0$ 1 2
$\alpha_{e}<0,\beta_{e}<0$ 1 2
From the Von Neumann and Krein theory, one can expect self-adjoint extensions of $iA_{0}$ only if the deficiency indices agree, that is, $d_{+}=d_{-}$. In such case, according to Theorem 2.3, there exist extensions that generate unitary dynamics. Several combinations of signs of $\alpha_{e}$ and $\beta_{e}$ along the edges of $\mathcal{G}$ may lead to equal deficiency indices. For instance, in case the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we would get deficiency indices $d_{+}$ and $d_{-}$ equal to $33k+$ 3. Thus, in this case, we know that there must be a $9(k+1)^{2}$ family of unitary groups for the case of the looping-edge graph, each one having its own representation.
4. The Airy operator on a looping-edge graph
As mentioned in the previous section, unitary dynamics are expected only in particular cases where the deficiency indices match for the free Airy operator. In the following subsections, we will introduce different approaches on a looping-edge graph $\mathcal{G}$ based on whether the deficiency indices coincide or not. We start by considering a ”pair-wise” reference frame on $\mathcal{G}$ that generates unitary dynamics, (see Theorem 4.3 below).
4.1. Boundary system for even amount of half-lines
Let us first consider $\mathcal{G}$ consisting of an even amount $2k$ of half-lines attached to the vertex in which exactly half of them have a leading coefficient $\alpha_{e}>0$. More precisely, we consider $\mathcal{G}$ defined trough $V=\{L,-L\}$ and $E=\{[-L,L]\}\cup E_{+}\cup E_{-}$ where $|E_{-}|=|E_{+}|=k\in\mathbb{Z}^{+}$ and $e\in E_{\pm}$ if and only if ±$\pm\alpha_{e}>0$.
Our objective in this section is to provide a first description of the boundary values of $A_{0}^{*}=-A_{0}$ ($D(A_{0}^{*})=H^{3}(\mathcal{G})$), we proceed as in Section 2 of [44]. For an alternative description see also Lemma 4.8 in Section 4.3.
Definition 4.1.
We define in the graph of $A_{0}^{*}=-A_{0}$,
$G(A_{0}^{*})=\{(\textbf{U},A_{0}^{*}\textbf{U})\mid\textbf{U}\in D(A_{0}^{*})\}$
the standard anti-symmetric form $\Omega:G(A_{0}^{*})\times G(A_{0}^{*})\to\mathbb{C},$ defined by
[…]
with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] being the $L^{2}(\mathcal{G})$-inner product in (3.5).
Proposition 4.1.
For any $\textbf{U}=(\phi,\psi_{1}\dots,\psi_{N})$ and $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$ in $D(A_{0}^{*})$ we have
$\begin{split}&[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\\ &\hskip 14.22636pt\left(\begin{pmatrix}\beta_{0}&0&\alpha_{0}\\ 0&-\alpha_{0}&0\\ \alpha_{0}&0&0\end{pmatrix}\begin{pmatrix}\phi(-L)\\ \phi^{\prime}(-L)\\ \phi^{\prime\prime}(-L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}(-L)\\ \tilde{\phi}^{\prime}(-L)\\ \tilde{\phi}^{\prime\prime}(-L)\end{pmatrix}\right)-\left(\begin{pmatrix}\beta_{0}&0&\alpha_{0}\\ 0&-\alpha_{0}&0\\ \alpha_{0}&0&0\end{pmatrix}\begin{pmatrix}\phi(L)\\ \phi^{\prime}(L)\\ \phi^{\prime\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}(L)\\ \tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime\prime}(L)\end{pmatrix}\right)\\ &\hskip 14.22636pt+\sum_{j=1}^{N}\left(\begin{pmatrix}\beta_{j}&0&\alpha_{j}\\ 0&-\alpha_{j}&0\\ \alpha_{j}&0&0\end{pmatrix}\begin{pmatrix}\psi_{j}(L)\\ \psi_{j}^{\prime}(L)\\ \psi_{j}^{\prime\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\psi}_{j}(L)\\ \tilde{\psi}_{j}^{\prime}(L)\\ \tilde{\psi}_{j}^{\prime\prime}(L)\end{pmatrix}\right)\\ &\hskip 14.22636pt=\left(B_{0}\partial\phi(-L)\mid\partial\tilde{\phi}(-L)\right)-\left(B_{0}\partial\phi(L)\mid\partial\tilde{\phi}(L)\right)+\sum_{j=1}^{N}\left(B_{j}\partial\psi_{j}(L)\mid\partial\tilde{\psi}_{j}(L)\right).\end{split}$ (4.2)
where
$B_{j}:=\begin{pmatrix}\beta_{j}&0&\alpha_{j}\\ 0&-\alpha_{j}&0\\ \alpha_{j}&0&0\end{pmatrix},\ j=0,\dots,N,$ (4.3)
[…]
Proof.
The proof follows by integration by parts. ∎
Next, we define the boundary-subspaces $\mathcal{B}_{\pm}$ as
+=$\mathcal{B}_{+}=\Big{\{}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big{\}}$ (4.4)
and
=$\mathcal{B}_{-}=\Big{\{}\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big{\}}.$ (4.5)
In this form, we have the following definition.
Definition 4.2.
Consider the spaces $\mathcal{B}_{+}\cong\mathcal{B}_{-}\cong\mathbb{C}^{3(k+1)}$ in (4.4)-(4.5) and endow these spaces with the non-degenerated indefinite inner products
$\left(x\mid y\right)_{\mathcal{B}_{+}}:=\left(x\mid y\right)_{+}=\left(B_{+}x\mid y\right)\mbox{ and }\ \left(x\mid y\right)_{\mathcal{B}_{-}}:=\left(x\mid y\right)_{-}:=\left(B_{-}x\mid y\right),$ (4.6)
[…]
$B_{+}:=\begin{pmatrix}B_{0}\\ &B_{1}\\ &&B_{3}\\ &&&\ddots\\ &&&&B_{2k-1}\end{pmatrix}\mbox{ and }B_{-}:=\begin{pmatrix}B_{0}\\ &-B_{2}\\ &&-B_{4}\\ &&&\ddots\\ &&&&-B_{2k}\end{pmatrix}.$
$(\mathcal{B}_{\pm},\left(\cdot\mid\cdot\right)_{\pm})$ are Krein spaces.
Under these new notations, Proposition 4.1 can be rewritten in a compact form as
$[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}_{odd}(L)\end{pmatrix}\right)_{+}-\left(\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \partial\tilde{\psi}_{even}(L)\end{pmatrix}\right)_{-}.$ (4.7)
Definition 4.3.
We define the function
$\begin{split}F:&\ G(A_{0}^{*})\to\mathcal{B}_{+}\oplus\mathcal{B}_{-}\\ &(\textbf{U},A_{0}^{*}\textbf{U})\mapsto\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix},\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}\right)\end{split}$ (4.8)
and the sesquilinear form
$\begin{split}w:&\ (\mathcal{B}_{+}\oplus\mathcal{B}_{-})\times(\mathcal{B}_{+}\oplus\mathcal{B}_{-})\to\mathbb{C}\\ &((x,y),(\tilde{x},\tilde{y}))\mapsto\left(x\mid\tilde{x}\right)_{+}-\left(y\mid\tilde{y}\right)_{-}.\end{split}$ (4.9)
Note from (4.1) and (4.2) that
$\Omega((\textbf{U},A_{0}^{*}\textbf{U}),(\textbf{V},A_{0}^{*}\textbf{V}))=w(F(\textbf{U},A_{0}^{*}\textbf{U}),F(\textbf{V},A_{0}^{*}\textbf{V})).$ (4.10)
By [44] we can now characterize skew-self-adjoint extensions $A$ of $A_{0}$ i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$. Note that obviously $G(A)=\{(x,Ax):x\in D(A)\}\subset G(A_{0}^{*})$
Proposition 4.2.
An extension $A$ of $A_{0}$ is skew-self-adjoint if and only if there exists a $w$-self-orthogonal subspace $X\subset\mathcal{B}_{+}\oplus\mathcal{B}_{-}$ such that $G(A)=F^{-1}(X)$.
Proof.
It follows immediately from Definitions 2.1 and 4.3, relations (4.7) and (4.10), and from [44]. ∎
Hence, $w$-self-orthogonal subspaces $X$ parametrize the skew-self-adjoint extensions $A$ of $A_{0}$. A more explicit description of these objects is given next.
4.1.1. Extensions generating unitary dynamics
Theorem 4.3.
Let $X\subset\mathcal{B}_{+}\oplus\mathcal{B}_{-}$. $X$ is $w$-self-orthogonal if and only if there is a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator $\mathcal{L}$ such that $X=G(\mathcal{L})$.
Proof.
We follow the same ideas of [38]. For sake of completeness we carry on with the details. Let $X$ be $w$-self-orthogonal and let $(x,y)\in X$. By definition we have $w((x,y),(x,y))=0$, that is,
$(x\mid x)_{+}=(y\mid y)_{-}.$ (4.11)
Since $X$ is a subspace, by setting $x=0$ we get $0=(y\mid y)_{-}$. Since $(\cdot\mid\cdot)_{-}$ is non-degenerated it must be the case $y=0$. By linearity we can conclude that $X$ is the graph of a linear operator $\mathcal{L}:\mathcal{B}_{+}\to\mathcal{B}_{-}$. Also note that if $y=0$ it also implies $x=0$; which means $\mathcal{L}$ is injective.
To see $\mathcal{L}$ is densely defined, let $z\in D(\mathcal{L})^{\perp}$ and $x\in D(\mathcal{L})$. We have
$0=(z\mid x)=(B_{+}B_{+}^{-1}z\mid x)=(B_{+}^{-1}z\mid x)_{+}=w((B_{+}^{-1}z,0),(x,\mathcal{L}x)).$ (4.12)
Since $x\in D(\mathcal{L})$ is arbitrary, we conclude $(B_{+}^{-1}z,0)\in X^{\perp_{w}}=X=G(\mathcal{L})$. Since $\mathcal{L}$ is injective, it must be the case $B^{-1}_{+}z=0$. The latter implies $z=0$. A symmetric argument shows the image $R(\mathcal{L})$ is dense.
Let us now see #$\mathcal{L}^{\#}=L^{-1}$. First note for $x\in D(\mathcal{L})$ and $y\in R(\mathcal{L})$, from (4.11), we have
$(\mathcal{L}x\mid y)_{-}=(x\mid\mathcal{L}^{-1}y)_{+}.$
The latter implies $R(\mathcal{L})\subset D(\mathcal{L}^{\#})$ with #$\mathcal{L}^{\#}y=\mathcal{L}^{-1}y$; i.e., $\mathcal{L}^{-1}\subset\mathcal{L}^{\#}$. On the other hand, if $y\in D(\mathcal{L}^{\#})$ with #$\mathcal{L}^{\#}y=x$ we have
$(\mathcal{L}z\mid y)_{-}=(z\mid x)_{+}\ \ \mbox{for all }z\in D(\mathcal{L}).$
Note the latter is equivalent to say
$w((z,\mathcal{L}z),(x,y))=0\ \mbox{ for all }z\in D(\mathcal{L}),$
which implies $(x,y)\in X^{\perp_{w}}=X=G(\mathcal{L})$. Thus, $(y,x)\in G(\mathcal{L}^{-1})$. We conclude $\mathcal{L}^{-1}=\mathcal{L}^{\#}$; proving $\mathcal{L}$ is unitary.
Conversely, let $\mathcal{L}$ be a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator with $X=G(\mathcal{L})$. Let $x\in D(\mathcal{L})$. For any $z\in D(\mathcal{L})$ we have $(\mathcal{L}x\mid\mathcal{L}z)_{-}=(x\mid z)_{+}$. That is, $w((x,\mathcal{L}x),(z,\mathcal{L}z))=0$ for any $z\in D(\mathcal{L})$. We conclude $(x,\mathcal{L}x)\in X^{\perp_{w}}$, i.e., $X\subset X^{\perp_{w}}$. On the other hand, if $(x,y)\in X^{\perp_{w}}$ we have $w((x,y),(z,\mathcal{L}z))=0$ for all $z\in D(\mathcal{L})$, i.e., $(\mathcal{L}z\mid y)_{-}=(z\mid x)_{+}$ for all $z\in D(\mathcal{L})$. By definition of $\mathcal{L}^{\#}$ the latter implies $y\in D(\mathcal{L}^{\#})=D(\mathcal{L}^{-1})$ with $\mathcal{L}^{-1}y=\mathcal{L}^{\#}y=x$. We conclude $(x,y)\in G(\mathcal{L})=X$. Hence $X=X^{\perp_{w}}$ as desired. ∎
4.1.2. Examples
With the first example below we intend to resemble the so-called $\delta$ interactions on a looping-edge graph $\mathcal{G}$ in a similar way as done in the case of metric star-shaped graphs (see, for instance, Theorem 3.6 in [8]).
Example 1 ($\delta$-type interactions on a looping-edge graph).
Let us first consider $k=1$, i.e., two half-lines attached to a common vertex. Assume $-\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $-\beta_{2}=\beta_{1}=\beta_{0}$. In particular,
$A_{0}=\left(\alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ \alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ -\alpha_{0}\partial_{x}^{3}-\beta_{0}\partial_{x}\right),$ (4.13)
and $B_{+}=B_{-}$ in Definition 4.2.
For parameters $z,m\in\mathbb{R}$ consider the operator
$\mathcal{L}:=\begin{pmatrix}1&0&0&0&0&0\\ z&1&0&0&0&0\\ \frac{z^{2}}{2}&z&1&m&0&0\\ 0&0&0&1&0&0\\ 0&0&0&z&1&0\\ -m&0&0&\frac{z^{2}}{2}&z&1\end{pmatrix}.$ (4.14)
Since #$\mathcal{L}^{\#}=B_{+}^{-1}\mathcal{L}^{*}B_{-}=\mathcal{L}^{-1}$, it can be seen that $\mathcal{L}$ defines a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator. According to Proposition 4.2 and Theorem 4.3, $\mathcal{L}$ defines a skew-self-adjoint extension $A_{\mathcal{L}}$ of $A_{0}$ defined via the relation $F(G(A_{\mathcal{L}}))=G(\mathcal{L})$. That is, an extension with domain $D(A_{\mathcal{L}})$ consisting of the functions $\textbf{V}=(\phi,\psi_{1},\psi_{2})\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(L)=\phi(-L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=z\phi(-L),\ \\ &\psi^{\prime}_{2}(L)-\psi^{\prime}_{1}(L)=z\psi_{1}(L),\ \phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=\frac{z^{2}}{2}\phi(-L)+z\phi^{\prime}(-L)+m\psi_{1}(L)\\ &\mbox{and }\psi_{2}^{\prime\prime}(L)-\psi_{1}^{\prime\prime}(L)=\frac{z^{2}}{2}\psi_{1}(L)+z\psi_{1}^{\prime}(L)-m\phi(-L).\end{split}$ (4.15)
Note in particular that if $\textbf{V}\in D(A_{\mathcal{L}})$ is so that $\phi(-L)=\psi_{1}(L)$ we may conclude from (4.15) that
$\begin{split}&\phi(-L)=\phi(L)=\psi_{1}(L)=\psi_{2}(L),\\ &\left(\phi^{\prime}(L)+\psi^{\prime}_{2}(L)\right)-\left(\phi^{\prime}(-L)+\psi^{\prime}_{1}(L)\right)=2z\phi(-L).\end{split}$ (4.16)
$-L$
$L$
$\phi$
$\psi_{1}$
$\psi_{2}$
$-L$
$L$
$\phi$
$\psi_{1}$
$\psi_{2}$
Figure 2. Examples of profiles satisfying (4.15) (on the left) and (4.16) (on the right).
More generally, consider $k\in\mathbb{Z}^{+}$ arbitrary (i.e. $2k$ half-lines attached to the common vertex $\nu=L$) and a parameter $z\in\mathbb{R}$. Under the analogous assumption $-\alpha_{2j}=\alpha_{2j-1}=\alpha_{0}$, $-\beta_{2j}=\beta_{2j-1}=\beta_{0}$ for all $j\in\{1,\dots,k\}$, so we obtain in particular for $k=2$
$A_{0}=\left(\alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ \alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ -\alpha_{0}\partial_{x}^{3}-\beta_{0}\partial_{x},\ \alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ -\alpha_{0}\partial_{x}^{3}-\beta_{0}\partial_{x}\right).$ (4.17)
Define the following operator with diagonal matrix of blocks
$\mathcal{L_{\delta_{Z}}}:=\begin{pmatrix}\delta_{z}\\ &\delta_{z}\\ &&\ddots\\ &&&\delta_{z}\end{pmatrix}\mbox{ where }\delta_{z}:=\begin{pmatrix}1&0&0\\ z&1&0\\ \frac{z^{2}}{2}&z&1\end{pmatrix}.$ (4.18)
Note that each block $\delta_{z}$ is such that $B_{0}^{-1}\delta_{z}^{*}B_{0}=\delta_{z}^{-1}$ for any $z\in\mathbb{R}$ ($B_{0}$ in (4.2)) and since $\mathcal{L}_{\delta_{z}}^{\#}$ is a diagonal matrix of blocks of the form $B_{0}^{-1}\delta_{z}^{*}B_{0}$ it can be seen that $\mathcal{L}_{\delta_{z}}^{\#}=\mathcal{L}_{\delta_{z}}^{-1}$. From there it is easy to see that $\mathcal{L}_{\delta_{z}}$ is a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator. From Proposition 4.2 and Theorem 4.3, $\mathcal{L}_{\delta_{z}}$ defines a skew-self-adjoint extension $A_{\mathcal{L}_{\delta_{z}}}$ of $A_{0}$ with domain
$D_{\delta_{z}}:=\left\{\textbf{V}\in H^{3}(\mathcal{G})\ \mid\delta_{z}\partial\phi(-L)=\partial\phi(L)\mbox{ and }\delta_{z}\partial\psi_{2j-1}(L)=\partial\psi_{2j}(L),\ j=1,\dots,k.\right\}$ (4.19)
Note from (4.19) we obtained ”pair-wise” the conditions in (4.15). More precisely, abusing notation to identify $\psi_{-1}$ with $\phi$ at $-L$ and $\psi_{0}$ with $\phi$ at $L$ we have for any $j=0,\dots,k$ that
$\begin{split}&\psi_{2j}(L)=\psi_{2j-1}(L),\ \psi_{2j}^{\prime}(L)-\psi_{2j-1}^{\prime}(L)=z\psi_{2j}(L)\mbox{ and}\\ &\psi_{2j}^{\prime\prime}(L)-\psi_{2j-1}^{\prime\prime}(L)=z\psi_{2j-1}^{\prime}(L)+\frac{z^{2}}{2}\psi_{2j-1}(L).\end{split}$ (4.20)
Moreover, if in particular $\textbf{V}\in D_{\delta_{z}}$ is such that $\phi(-L)=\psi_{2j-1}(L)$ for all $j=1,\dots,k$ then we conclude from (4.19) that
$\phi(-L)=\phi(L)=\psi_{1}(L)=\cdots=\psi_{2k}(L)\mbox{ and }\sum_{j=0}^{k}\psi^{\prime}_{2j-1}(L)-\sum_{j=0}^{k}\psi^{\prime}_{2j}(L)=(1+k)z\phi(-L).$ (4.21)
For the following examples we assume $|E_{+}|=|E_{-}|=1$. From the deficiency indices in Section 3 and the Von Neumann and Krein theory we know the skew-self-adjoint extensions of $A_{0}$ are parametrized by $366^{2}=$ 36 parameters. We exhibit a $2$ and a $4$-parameter independent families of unitary operators that maintain the looping-edge structure (i.e. the loop condition $\phi(-L)=\phi(L)$) and are not necessarily interacting pair-wise.
Example 2.
Consider again $k=1$ and $B_{+}=B_{-}$. It can be seen for $m_{1},m_{2}\in\mathbb{R}$ that
$\mathcal{L}=\begin{pmatrix}1&0&0&0&0&0\\ 0&0&0&m_{1}&1&0\\ 0&0&1&m_{2}&0&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ -m_{2}&0&0&\frac{m_{1}^{2}}{2}&m_{1}&1\end{pmatrix}$ (4.22)
is a unitary operator and therefore that it defines a skew-self-adjoint extension $A_{\mathcal{L}}$ with domain $D(A_{\mathcal{L}})$ made of the functions $v=(\phi,\psi_{1},\psi_{2})\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(-L)=\psi^{\prime}_{2}(L),\ \phi^{\prime}(-L)-\psi^{\prime}_{1}(L)=m_{1}\psi_{1}(L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=m_{2}\psi_{1}(L)\ \mbox{and }\psi_{2}^{\prime\prime}(L)-\psi_{2}^{\prime\prime}(-L)=\frac{m_{1}^{2}}{2}\psi_{1}(L)+m_{1}\psi_{1}^{\prime}(L)-m_{2}\phi(-L).\end{split}$
Example 3.
Consider again $k=1$, $-\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $-\beta_{2}=\beta_{1}=\beta_{0}$. It can be seen for $m_{1},\dots,m_{4}\in\mathbb{R}$ that
$\mathcal{L}=\begin{pmatrix}1&0&0&0&0&0\\ m_{1}&0&0&m_{2}&1&0\\ \frac{m_{1}^{2}+1}{2}&1&1&m_{3}&m_{1}&0\\ 0&0&0&1&0&0\\ 1&1&0&m_{4}&0&0\\ m_{4}-m_{3}+m_{1}m_{2}&m_{4}&0&\frac{m_{2}^{2}+m_{4}^{2}}{2}&m_{2}&1\end{pmatrix}$ (4.23)
defines a skew-self-adjoint extension $A_{\mathcal{L}}$ with domain $D(A_{\mathcal{L}})$ made of $v\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(L)=\phi(-L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=m_{1}\phi(-L)+m_{2}\psi_{1}(L),\\ &\psi_{2}^{\prime}(L)-\phi^{\prime}(-L)=m_{4}\psi_{1}(L)+\phi(-L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=\frac{m_{1}^{2}+1}{2}\phi(-L)+m_{3}\psi_{1}(L)+\phi^{\prime}(-L)+m_{1}\psi_{1}^{\prime}(L),\mbox{and}\\ &\psi_{2}^{\prime\prime}(L)-\psi_{1}^{\prime\prime}(L)=(m_{4}-m_{3}+m_{1}m_{2})\phi(-L)+\frac{m_{2}^{2}+m_{4}^{2}}{2}\psi_{1}(L)+m_{4}\phi^{\prime}(L)+m_{2}\psi_{1}^{\prime}(L).\end{split}$
4.2. Boundary system for arbitrary amount of half-lines
For this second part we no longer seek for unitary dynamics. We feel confortable allowing the dynamics to be generated by a $C_{0}$-semigroup of contractions. To do so, we consider $\mathcal{G}$ consisting of finite amount $N$ of half-lines attached to the vertex $L$ with no restrictions on the signs of the coefficients of the Hamiltonian. We still treat $-L$ as a terminal vertex but we will be emphatic in whether the examples to be presented maintain or not the looping-edge structure. It is worth mentioning that most of the following results also apply to tadpole-like graphs.
Definition 4.4.
Consider the spaces $\mathcal{C}_{1}\cong\mathcal{C}_{0}\cong\mathbb{C}^{3(N+1)}$ and endow these spaces with the non-degenerated indefinite inner products
$\left(x\mid y\right)_{1}:=\left(C_{1}x\mid y\right)\mbox{ and }\ \left(x\mid y\right)_{0}:=\left(C_{0}x\mid y\right),$ (4.24)
where
$C_{1}:=\begin{pmatrix}B_{0}\\ &B_{1}\\ &&B_{2}\\ &&&\ddots\\ &&&&B_{N}\end{pmatrix}\mbox{ and }C_{0}:=\frac{1}{N+1}\begin{pmatrix}B_{0}\\ &B_{0}\\ &&B_{0}\\ &&&\ddots\\ &&&&B_{0}\end{pmatrix}.$
$(\mathcal{C}_{1},\left(\cdot\mid\cdot\right)_{1})$ and $(\mathcal{C}_{0},\left(\cdot\mid\cdot\right)_{0})$ are Krein spaces.
Under these new notations, Proposition 4.1 can be rewritten as
$[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}-\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}.$ (4.25)
Definition 4.5.
We consider the function
$\begin{split}\mathcal{F}:&\ G(A_{0}^{*})\to\mathcal{C}_{1}\oplus\mathcal{C}_{0}\\ &(\textbf{U},A_{0}^{*}\textbf{U})\mapsto\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix},\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)\end{split}$ (4.26)
4.2.1. Extensions generating contractile dynamics
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be a densely defined operator. Consider an extension $A_{\mathcal{L}}$ of $A_{0}$ defined by the relation $G(A_{\mathcal{L}})=\mathcal{F}^{-1}(G(\mathcal{L}))$. In other words,
$D(A_{\mathcal{L}}):=\left\{\textbf{U}\in D(A_{0}^{*})\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\in D(\mathcal{L})\mbox{ and }\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right\}.$ (4.27)
acting as $A_{\mathcal{L}}\textbf{U}:=-A_{0}^{*}\textbf{U}$.
Let us determine the adjoint of operator of $A_{\mathcal{L}}$.
Proposition 4.4.
For $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ densely defined we have
$D(A_{\mathcal{L}}^{*}):=\left\{(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{0}^{*})\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ and }\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right\}$ (4.28)
acting as $A_{\mathcal{L}}^{*}\textbf{U}=A_{0}^{*}\textbf{U}.$
Proof.
Let $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{\mathcal{L}})$, $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})\in D(A_{0}^{*})$. From (4.7) and the definition of $A_{\mathcal{L}}$ we have
$\begin{split}[A_{\mathcal{L}}\textbf{U},\textbf{V}]&=[\textbf{U},A_{0}^{*}\textbf{V}]+\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}\\ &=[\textbf{U},A_{0}^{*}\textbf{V}]+\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}.\end{split}$ (4.29)
Therefore, $\textbf{V}\in D(A_{\mathcal{L}}^{*})$ if and only if for all $\textbf{U}\in D(A_{\mathcal{L}})$ it follows
$\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}.$ (4.30)
Note from the definition of $A_{\mathcal{L}}$ that $\mathcal{F}(G(A_{\mathcal{L}}))=G(\mathcal{L})$; which means
$D(\mathcal{L})=\left\{\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\textbf{U}\in D(A_{\mathcal{L}})\right\}.$ (4.31)
We conclude (4.30) is equivalent to
#$\mathcal{L}^{\#}\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}=\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}.$
From Proposition 4.4 we see that knowing the particular expression defining $\mathcal{L}^{\#}$ is important to determine explicitly the domain of $A_{\mathcal{L}}^{*}$. Note in case $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ is given by a $3(N+1)\times 3(N+1)$ matrix, the $(\mathcal{C}_{0},\mathcal{C}_{1})$-adjoint $\mathcal{L}^{\#}$ is given by the expression
#$\mathcal{L}^{\#}=C_{1}^{-1}\mathcal{L}^{*}C_{0}.$ (4.32)
In fact, for $y\in\mathcal{C}_{0}$ the definition of $\mathcal{L}^{\#}$ imply that for all $x\in\mathcal{C}_{1}$ we have
$\left(C_{0}\mathcal{L}x\mid y\right)=\left(\mathcal{L}x\mid y\right)_{0}=\left(x\mid\mathcal{L}^{\#}y\right)_{1}=\left(C_{1}x\mid\mathcal{L}^{\#}y\right)_{0}.$ (4.33)
Since $C_{1}$ is real symmetric we obtain that for all $x\in\mathcal{C}_{1}$
$\left(x\mid\ \left(\mathcal{L}^{*}C_{0}-C_{1}\mathcal{L}^{\#}\right)y\right)=0.$ (4.34)
Since $y$ is arbitrary, it follows
$\mathcal{L}^{*}C_{0}=C_{1}\mathcal{L}^{\#},$ (4.35)
which implies (4.32)
In the following two propositions we characterize when both, $A_{\mathcal{L}}$ and $A_{\mathcal{L}}^{*}$, are dissipative operators. We recall that dissipative operators on Hilbert spaces were briefly described on subsection 2.2.
Proposition 4.5.
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$. $A_{\mathcal{L}}$ is dissipative if and only if $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction.
Proof.
Let $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{\mathcal{L}})$. From (4.7) and the definition of $A_{\mathcal{L}}$ we have
$\begin{split}2\mbox{Re}[A_{\mathcal{L}}\textbf{U},\textbf{U}]&=[A_{\mathcal{L}}\textbf{U},\textbf{U}]+\overline{[A_{\mathcal{L}}\textbf{U},\textbf{U}]}=[A_{\mathcal{L}}\textbf{U},\textbf{U}]+[\textbf{U},A_{\mathcal{L}}\textbf{U}]=-[A_{0}^{*}\textbf{U},\textbf{U}]-[\textbf{U},A_{0}^{*}\textbf{U}]\\ &=\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{1}.\end{split}$ (4.36)
Hence, $\mbox{Re}[A_{\mathcal{L}}\textbf{U},\textbf{U}]\leqslant 0$ if and only if for all $\textbf{U}\in D(A_{\mathcal{L}})$ it follows
$\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{0}\leqslant\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{1}.$ (4.37)
From (4.31) we conclude that (4.37) is equivalent to $\mathcal{L}$ being a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction. ∎
Proposition 4.6.
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be densely defined. $A_{\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction.
Proof.
Let $\textbf{V}\in D(A_{\mathcal{L}}^{*})$. From (4.7) we have
$\begin{split}2\mbox{Re}[A_{\mathcal{L}}^{*}\textbf{V},\textbf{V}]&=[A_{0}^{*}\textbf{V},\textbf{V}]+[\textbf{V},A_{0}^{*}\textbf{V}]\\ &=\left(\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{1}-\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{0}.\end{split}$ (4.38)
Hence, $\mbox{Re}[A_{\mathcal{L}}^{*}\textbf{V},\textbf{V}]\leqslant 0$ if and only if
(#$\left(\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{1}\leqslant\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{0}\ \mbox{ for all }\textbf{V}\in D(A_{\mathcal{L}}^{*}).$ (4.39)
From the definition of $D(A_{\mathcal{L}}^{*})$ we conclude (4.39) is equivalent to $\mathcal{L}^{\#}$ being a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction. ∎
From the previous two propositions above and the Lummer-Phillips Theorem (see Theorem 2.2) we can obtain the following characterization of extensions generating $C_{0}$-semigroups of contractions.
Theorem 4.7.
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be densely defined. $A_{\mathcal{L}}$ is the generator of a contraction semigroup if and only if $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction and $\mathcal{L}^{\#}$ satisfies (4.39).
Proof.
Suppose $A_{\mathcal{L}}$ generates a semigroup $\{T(t)\}_{t\geqslant 0}$ of contractions. By Proposition 2.1, $A_{\mathcal{L}}$ is dissipative. Since $A_{\mathcal{L}}^{*}$ generates the semigroup $\{T(t)^{*}\}_{t\geqslant 0}$ which is also a contraction semigroup, from Proposition 2.1 we conclude $A_{\mathcal{L}}^{*}$ is dissipative. From Propositions 4.5 and 4.6 we conclude that $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction and $\mathcal{L}^{\#}$ satisfies (4.39). The converse follows form Propositions 4.5 and 4.6 and Theorem 2.2. ∎
4.2.2. Examples
Let us show an example an extension $A_{\mathcal{L}}$ generating semigroups of contractions. We intent to present an operator whose first row preserves the looping-edge structure ($\phi(L)=\phi(-L)$) with continuity at the vertex and still keeping several parameters. For simplicity we consider a tadpole graph ($N=1$) with $\alpha_{0}=\alpha_{1}=\beta_{0}=\beta_{1}$.
Example 4 ($\delta$ type interactions on a tadpole graph).
Let $m_{1},\dots,m_{4}\in\mathbb{R}$ and let
$M:=\begin{pmatrix}1&0&0&0&0&0\\ \frac{m_{2}+m_{3}}{2}&1&0&1&0&0\\ m_{1}&m_{2}+m_{3}&2&m_{2}&0&0\\ 0&0&0&1&0&0\\ \frac{m_{2}+m_{3}}{2}&1&0&1&0&0\\ m_{3}&2&0&m_{4}&0&2\end{pmatrix}.$ (4.40)
According to (2.4), $M$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction if for all $x\in\mathbb{C}^{6}$
$\left(C_{0}Mx\mid Mx\right)-\left(C_{1}x\mid x\right)\leqslant 0,$ (4.41)
i.e., if the matrix $M^{*}C_{0}M-C_{1}$ is negative semidefinite.
It can be seen that it is enough to consider
$\alpha_{0}<0,\ m_{4}\leqslant\frac{3}{2}\mbox{ and }4m_{1}-(m_{2}^{2}+m_{3}^{2})\leqslant 2.$ (4.42)
Now, according to (4.32) we have
$M^{\#}=\begin{pmatrix}1&0&0&0&0&0\\ \frac{-m_{2}-m_{3}}{2}&\frac{1}{2}&0&-1&\frac{1}{2}&0\\ \frac{m_{1}-1}{2}&\frac{-m_{2}-m_{3}}{4}&\frac{1}{2}&\frac{m_{3}}{2}&\frac{-m_{2}-m_{3}}{4}&0\\ 0&0&0&1&0&0\\ 0&0&0&0&0&0\\ \frac{m_{2}}{2}&\frac{-1}{2}&0&\frac{m_{4}-1}{2}&\frac{-1}{2}&\frac{1}{2}\end{pmatrix}.$ (4.43)
In order to $M_{z}^{\#}$ be a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction we require for all $y\in\mathbb{C}^{6}$ that
$\left(C_{1}M^{\#}y\mid M^{\#}y\right)\leqslant\left(C_{0}y\mid y\right),$ (4.44)
i.e., we require the matrix $(M^{\#})^{*}C_{1}M^{\#}-C_{0}$ to be negative semidefinite The latter holds under the exact same condition (4.42).
The operator $M$ defines an operator $A_{M}$ acting as $A_{M}\textbf{U}=-A_{0}^{*}\textbf{U}$ with domain consisting of functions $(\phi,\psi)\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=\left(\frac{m_{2}+m_{3}+2}{2}\right)\phi(-L),\\ &\phi^{\prime\prime}(L)-2\psi^{\prime\prime}(L)=(m_{3}+m_{4})\psi(L)+2\phi^{\prime}(-L)\ \mbox{ and }\\ &\phi^{\prime\prime}(L)-2\phi^{\prime\prime}(-L)=(m_{1}+m_{2})\phi(-L)+(m_{2}+m_{3})\phi^{\prime}(L).\end{split}$ (4.45)
Note if you take the parameters satisfying $m_{4}=m_{1}=0$, $m_{3}=-2$ and denoting $z=m_{2}/2\in\mathbb{R}$ it follows from (4.45)
$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=z\phi(-L)\mbox{ and }\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)-\psi^{\prime\prime}(-L)=z\phi^{\prime}(-L)+(z-1)\phi(-L).\end{split}$ (4.46)
We conclude from Proposition 4.5, Proposition 4.6 and Theorem 4.7 that $A_{M}$ generates a semigroup of contractions.
4.3. Boundary system separating derivatives
A better description of the boundary behavior can be performed by separating the relation on the first derivatives exhibited by the matrices $B_{j}$ in (4.2). It can be attained by rewriting Proposition 4.1 in the following way. We recall the notation associated to U, $\vec{\textbf{{U}}}$, $\vec{\textbf{{U}}}^{\prime}$ and $\vec{\textbf{{U}}}^{\prime\prime}$ given in subsection 2.1.
Lemma 4.8.
For any $\textbf{U}=(\phi,\psi_{1}\dots,\psi_{N})$ and $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$ in $D(A_{0}^{*})$ we have
$\begin{split}&[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(D_{\beta}\vec{\textbf{{U}}}\mid\vec{\textbf{V}}\right)+\left(D_{\alpha}\vec{\textbf{{U}}}^{\prime\prime}\mid\vec{\textbf{V}}\right)+\left(D_{\alpha}\vec{\textbf{{U}}}\mid\vec{\textbf{V}}^{\prime\prime}\right)-\left(D_{\alpha}\vec{\textbf{{U}}}^{\prime}\mid\vec{\textbf{V}}^{\prime}\right),\end{split}$ (4.47)
where
$\begin{split}&D_{\alpha}:=\begin{pmatrix}\alpha_{0}&&&&\\ &-\alpha_{0}&&&&\\ &&\alpha_{1}&&\\ &&&\ddots&\\ &&&&\alpha_{N}\end{pmatrix}\mbox{ and }\ D_{\beta}:=\begin{pmatrix}\beta_{0}&&&&\\ &-\beta_{0}&&&&\\ &&\beta_{1}&&\\ &&&\ddots&\\ &&&&\beta_{N}\end{pmatrix}.\end{split}$ (4.48)
Definition 4.6.
Consider the spaces $\mathcal{K}_{+}\cong\mathcal{K}_{-}\cong\mathbb{C}^{N+1}$ and endow them with the weighted products
$\left(x\mid y\right)_{\mathcal{K}_{+}}:=\left(D^{+}_{\alpha}x\mid y\right)\mbox{ and }\left(x\mid y\right)_{\mathcal{K}_{-}}:=\left(D_{\alpha}^{-}x\mid y\right),$ (4.49)
where
$\begin{split}&D_{\alpha}^{+}:=\begin{pmatrix}\alpha_{0}&&&&\\ &\alpha_{1}&&&&\\ &&\alpha_{2}&&\\ &&&\ddots&\\ &&&&\alpha_{N}\end{pmatrix}\mbox{ and }\ D_{\alpha}^{-}:=\frac{1}{N+1}\begin{pmatrix}\alpha_{0}&&&&\\ &\alpha_{0}&&&&\\ &&\alpha_{0}&&\\ &&&\ddots&\\ &&&&\alpha_{0}\end{pmatrix}.\end{split}$
Then $(\mathcal{K}_{\pm},(\cdot\mid\cdot)_{\mathcal{K}_{\pm}})$ are Krein spaces.
Note the matrices $D_{\alpha}^{-}$ and $D_{\alpha}^{+}$ are simply splitting the matrix $D_{\alpha}$ into two matrices, one dealing with the coefficient of $\phi(L)$ and another dealing with the rest of the coefficients. In fact,
$\left(D_{\alpha}\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}^{\prime}\right)=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.$ (4.50)
From (4.47) and (4.50) we can rewrite
$\begin{split}[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]&=\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\mid\vec{\textbf{V}}\right)+\left(\vec{\textbf{U}}\mid\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)-\left(D_{\alpha}\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}^{\prime}\right)\\ &=\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\mid\vec{\textbf{V}}\right)+\left(\vec{\textbf{U}}\mid\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)\\ &\hskip 28.45274pt-\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}+\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.51)
4.3.1. Extensions generating contractile dynamics via splitting derivatives
Let us introduce extensions of $A_{0}$ that are parameterized by two items: a linear subspace $Y$ that characterize the interaction between the function and its second derivatives and a linear operation $\mathcal{L}$ that governs how the first derivatives relate each other. Different than in the case of subsection 4.2.1, we cannot obtain mixed expression involving the function and its first derivatives with these extensions $A_{Y,\mathcal{L}}$ defined next. One possible application of the results below, can be to investigate the existence and stability of solitary waves for the KdV model (1.6) on looping edge graphs in terms of the coefficients of the equations on each edge (see [39]-section 3 and Theorem 4.2).
Definition 4.7.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and consider a closed subspace $Y\subset\mathbb{C}^{N+2}$. We define $A_{Y,\mathcal{L}}$ on the set
$\begin{split}D(A_{Y,\mathcal{L}}):=\left\{\textbf{U}\in D(A_{0}^{*})\mid\right.&\vec{\textbf{U}}\in Y,\ D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\in Y^{\bot}\mbox{ and }\\ &\hskip 8.53581pt\left.\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\right\},\end{split}$ (4.52)
acting as $A_{Y,\mathcal{L}}\textbf{U}=-A_{0}^{*}\textbf{U}$.
The following proposition describe explicitly the form of the adjoint operator $A_{Y,\mathcal{L}}^{*}$ of $A_{Y,\mathcal{L}}$.
Proposition 4.9.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and consider a closed subspace $Y\subset\mathbb{C}^{N+2}$. We have that the adjoint operator $A_{Y,\mathcal{L}}^{*}$ of $A_{Y,\mathcal{L}}$ is defined on the domain
$\begin{split}D(A_{Y,\mathcal{L}}^{*})=\{\textbf{V}\in D(A_{0}^{*})\mid&\vec{\textbf{V}}\in Y,\ \frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\in Y^{\bot}\mbox{ and }\\ &\hskip 8.53581pt\left.\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ with }\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\right\}\end{split}$ (4.53)
acting as $A_{Y,\mathcal{L}}^{*}\textbf{V}=A_{0}^{*}\textbf{V}$.
Proof.
Let $\textbf{U}\in D(A_{Y,\mathcal{L}})$ and $\textbf{V}\in D(A_{0}^{*})$. We have
$\begin{split}[A_{Y,\mathcal{L}}\textbf{U},\textbf{V}]&=[-A_{0}^{*}\textbf{U},\textbf{V}]\\ &=[\textbf{U},A_{0}^{*}\textbf{V}]-\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\mid\vec{\textbf{V}}\right)-\left(\vec{\textbf{U}}\mid\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)\\ &\ +\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.54)
In order to V be in $D(A_{Y,\mathcal{L}}^{*})$ it must satisfy $[A_{Y,\mathcal{L}}\textbf{U},\textbf{V}]=[\textbf{U},A_{Y,\mathcal{L}}^{*}\textbf{V}]$ for all $\textbf{U}\in D(A_{Y,\mathcal{L}})$. Let $\textbf{U}\in D(A_{Y,\mathcal{L}})$ such that ${\textbf{U}}=c{\textbf{U}}^{\prime}=\vec{\textbf{0}}$. It follows from (4.54) that $\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}\mid\vec{\textbf{V}}\right)=0$. Since by definition of $A_{Y,\mathcal{L}}$ we have $D_{\alpha}\vec{\textbf{U}}^{\prime\prime}\in Y^{\bot}$, we conclude
$\vec{\textbf{V}}\in Y.$ (4.55)
Now, Let $x\in Y$ be arbitrary. There exists $\textbf{U}\in D(A_{Y,\mathcal{L}})$ such that $\vec{\textbf{U}}=x$ with $\vec{\textbf{U}}^{\prime}=0$. Since $D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}$ is in $Y^{\bot}$, from (4.54) we get
$\left(x,\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)=\left(\vec{\textbf{U}},\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)=\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}},\vec{\textbf{V}}\right)=0.$ (4.56)
We conclude
$\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\in Y^{\bot}$ (4.57)
From (4.54), (4.55) and (4.57) we obtain that for all $\textbf{U}\in D(A_{Y,\mathcal{L}})$ we must have
$\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}=\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.$ (4.58)
From definition of $D(A_{Y,\mathcal{L}})$, (4.58) is equivalent to have that for all $\textbf{U}\in D(A_{Y,\mathcal{L}})$
$\left(\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}.$ (4.59)
It is clear that
$\left\{\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}^{\prime}_{1}(L)\\ \vdots\\ \tilde{\psi}^{\prime}_{N}(L)\end{pmatrix}\mid\textbf{V}\in D(A_{Y,\mathcal{L}})\right\}\subset D(\mathcal{L}).$
Moreover, the equality holds since for any $x\in D(\mathcal{L})$ there exists $\textbf{V}\in D(A_{Y,\mathcal{L}})$ such that $x=\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}^{\prime}_{1}(L)\\ \vdots\\ \tilde{\psi}^{\prime}_{N}(L)\end{pmatrix}$. Therefore, (4.59) is equivalent to say
$\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ with }\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}.$ (4.60)
Proposition 4.10.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace. $A_{Y,\mathcal{L}}$ is dissipative if and only if $\mathcal{L}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction.
Proof.
Let $\textbf{U}\in D(A_{Y,\mathcal{L}})$. We have
$\begin{split}2\mbox{Re}[A_{Y,\mathcal{L}}\textbf{U},\textbf{U}]&=[A_{Y,\mathcal{L}}\textbf{U},\textbf{U}]+[\textbf{U},A_{Y,\mathcal{L}}\textbf{U}]=-[A_{0}^{*}\textbf{U},\textbf{U}]-[\textbf{U},A_{0}^{*}\textbf{U}]\\ &=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}_{N}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}\\ &=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.61)
Since $\textbf{U}\in D(A_{Y,\mathcal{L}})$ is arbitrary we conclude (arguing as in the proof of Proposition 4.9) that $\mbox{Re}[A_{Y,\mathcal{L}}\textbf{U},\textbf{U}]\leqslant 0$ if and only if $\mathcal{L}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction. ∎
The following two propositions characterize when the extensions $A_{Y,\mathcal{L}}$ are dissipative.
Proposition 4.11.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace. $A_{Y,\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{K}_{-},\mathcal{K}_{+})$-contraction.
Proof.
Let $\textbf{V}\in D(A_{Y,\mathcal{L}}^{*})$. We have
$\begin{split}2\mbox{Re}[A_{Y,\mathcal{L}}^{*}\textbf{V},\textbf{V}]&=[A_{0}^{*}\textbf{V},\textbf{V}]+[\textbf{V},A_{0}^{*}\textbf{V}]=\\ &=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}-\left(\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}.\end{split}$ (4.62)
It is clear that
$\left\{\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\textbf{V}\in D(A_{Y,\mathcal{L}}^{*})\right\}\subset D(\mathcal{L}^{\#}).$
Moreover, the equality holds since for any $y\in D(\mathcal{L}^{\#})$ there is $\textbf{V}\in D(A_{Y,\mathcal{L}}^{*})$ such that
$y=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}.$
From (4.62) we conclude $A_{Y,\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{K}_{-},\mathcal{K}_{+})$-contraction.
The following theorem is a direct consequence of the previous two propositions and can be proved as done in the proof of Theorem 4.7.
Theorem 4.12.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace. $A_{Y,\mathcal{L}}$ generates a contraction semigroup if and only if $\mathcal{L}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction and $\mathcal{L}^{\#}$ is a $(\mathcal{K}_{-},\mathcal{K}_{+})$-contraction.
4.3.2. Examples
Let us first describe the operators $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ that define contractions so that its respective adjoint $\mathcal{L}^{\#}$ is also a contraction. Let us assume for simplicity $N=1$ (i.e. a tadpole graph) and $\alpha_{0}=\alpha_{1}$.
Example 5.
For $m_{1},m_{2},m_{3}\in\mathbb{R}$ with $m_{3}\neq 0$, consider the operator
$\mathcal{L}:=\begin{pmatrix}m_{1}&m_{2}\\ m_{3}&\frac{-m_{1}m_{2}}{m_{3}}\end{pmatrix}.$
Note we have
#$\mathcal{L}^{\#}=\begin{pmatrix}\frac{m_{1}}{2}&\frac{m_{3}}{2}\\ \frac{m_{2}}{2}&\frac{-m_{1}m_{2}}{2m_{3}}\end{pmatrix}.$
It can be seen that both, $\mathcal{L}$ and $\mathcal{L}^{\#}$ are contractions under the conditions
$m_{3}^{2}+m_{1}^{2}\leqslant 2,\ m_{2}^{2}(m_{1}^{2}+1)\leqslant 2,\ m_{1}(m_{3}^{2}-m_{2}^{2})=0,\ m_{1}^{2}+m_{2}^{2}\leqslant 1\ \mbox{and }\frac{m_{1}^{2}m_{2}^{2}}{m_{3}^{2}}+m_{3}^{2}\leqslant 2.$ (4.63)
In case $m_{3}=m_{2}\neq 0$, (4.63) reduces to
$m_{2}^{2}+m_{1}^{2}\leqslant 2\mbox{ and }m_{2}^{2}(m_{1}^{2}+1)\leqslant 2.$
In particular, if we take $m_{2}=m_{1}=1$ you get that
$\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$
defines a contraction with $(\mathcal{K}_{-},\mathcal{K}_{+})$-adjoint also a contraction.
The next example pretends to characterize the extensions obtained via separating derivatives that maintain the looping-edge structure ($\phi(-L)=\phi(L)$) and include continuity at the vertex (when $z=1$ below).
Example 6.
Let $z\in\mathbb{R}$ be fixed and $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a linear operator. Consider the subspace $Y_{z}:=\mbox{span}\{(1,1,z)^{T}\}$. Let us first assume $z\neq 0$. We have that
$Y_{z}^{\perp}=\left\{(x,y,w)^{T}\in\mathbb{C}^{3}\mid w=\frac{-x-y}{z}\right\}.$
From the Definition 4.7 we have that the domain of $A_{Y_{z},\mathcal{L}}$ is made of functions $u\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L),\ \psi(L)=z\phi(L),\ z\left(\alpha_{1}\psi^{\prime\prime}(L)+\frac{\beta_{1}}{2}\psi(L)\right)=\alpha_{0}\left(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)\right)\\ &\mbox{and }\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}.\end{split}$ (4.64)
For instance, if we take $z=1$ and $\mathcal{L}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ we have that $A_{Y_{1},\mathcal{L}}$ with domain
$\begin{split}D(A_{Y_{1},\mathcal{L}})=\{\textbf{U}\in D(A_{0}^{*})\mid&\ \phi(L)=\phi(-L)=\psi(L),\ \phi^{\prime}(L)=\phi^{\prime}(-L),\\ &\ \psi^{\prime}(L)=0\mbox{ and }\alpha_{1}\psi^{\prime\prime}(L)+\frac{\beta_{1}}{2}\psi(L)=\alpha_{0}(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L))\}.\end{split}$ (4.65)
defines a contractile dynamic on the tadpole graph.
$-L$
$L$
$\psi$
$\phi$
Figure 3. Example of profile in (4.66)
On the other hand, assume $z=0$. In such case
$Y_{0}^{\perp}=\{(x,y,w)\in\mathbb{C}^{3}\mid x=-y\}.$
Therefore the domain of $A_{Y_{0},\mathcal{L}}$ is made of functions $\textbf{U}\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L),\ \psi(L)=0,\ \phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\ \mbox{and }\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}.\end{split}$
In particular if we take $\mathcal{L}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ we get that $A_{Y_{0},\mathcal{L}}$ with domain
$\begin{split}D(A_{Y_{0},\mathcal{L}})=\{\textbf{U}\in D(A_{0}^{*})\mid&\ \phi(L)=\phi(-L),\ \psi(L)=0,\ \phi^{\prime}(L)=\phi^{\prime}(-L)\mbox{ and }\phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\}.\end{split}$ (4.66)
defines an extension of $A_{0}$ generating a contractive dynamic on the tadpole.
5. The Schrödinger operator on looping-edge and $\mathcal{T}$-shaped graphs
In this section we employ the same techniques used in the case of the free Airy operator to characterize the extensions of the Schrödinger operator on a looping-edge and a graph determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering $-L$ as a terminal vertex. This strategy will allow us in the future to study the dynamics of NLS model (1.1) specially on $\mathcal{T}$-graphs.
Let us now consider the Schrödinger operator
$\mathcal{H}_{0}:\{u_{e}\}_{e\in E}\mapsto\left\{-\frac{d^{2}}{dx_{e}^{2}}u_{e}\right\}_{e\in E},$ (5.1)
seen as a densely defined operator on $L^{2}(\mathcal{G})$ with domain $D(\mathcal{H}_{0})$ defined on (1.3). Note we would have $D(\mathcal{H}_{0}^{*})=H^{2}(\mathcal{G})$.
5.1. Deficiency indices of the free Schrödinger operator
For the symmetric operator $\mathcal{H}_{0}$ in (5.1) we have that free solutions of $-\partial_{x}^{2}u_{e}\pm iu_{e}=0$ are given as linear combinations of complex exponentials of the form
$u_{e}=c_{1}e^{\frac{(1\pm i)}{\sqrt{2}}x_{e}}+c_{2}e^{-\frac{(1\pm i)}{\sqrt{2}}x_{e}}.$
We note that $e^{\pm\frac{(1\pm i)}{\sqrt{2}}x}$ belong to $L^{2}(-L,L)$. On the other hand, on a half-line $[L,\infty)$, it would be square integrable only if the real part of ±$\pm\frac{(1\pm i)}{\sqrt{2}}$ is negative. Along each edge $e=[L,\infty)$ we would therefore always have that the deficiency indices are ${d_{+}}_{e}={d_{-}}_{e}=1$.
In a looping-edge graph with $N$ half-lines attached at the vertex we would have for $\mathcal{H}_{0}$ that the deficiency indices $d_{\mp}=dim(Ker(\mathcal{H}_{0}\pm iI))$ satisfy $d_{+}=d_{-}=2+N$. From the von Neumann and Krein theory, one can always expect self-adjoint extensions of $\mathcal{H}_{0}$ and therefore unitary dynamics.
5.2. Boundary systems
In this subsection we adapt the ideas developed in the case of the Airy operator, for the Schrödinger operator posed on $\mathcal{G}$.
Definition 5.1.
We define in the graph of $\mathcal{H}_{0}^{*}$, $G(\mathcal{H}_{0}^{*})=\{(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})\mid\textbf{U}\in D(\mathcal{H}_{0}^{*})\}$, the standard symmetric form $\Lambda:G(\mathcal{H}_{0}^{*})\times G(\mathcal{H}_{0}^{*})\to\mathbb{C},$ defined by
[…]
Proposition 5.1.
For any $\textbf{U},\textbf{V}\in D(\mathcal{H}_{0}^{*})$ we have
$[\mathcal{H}_{0}^{*}\textbf{U},\textbf{V}]-[\textbf{U},\mathcal{H}_{0}^{*}\textbf{V}]=\left(P_{+}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)-\left(P_{-}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)$ (5.3)
where
$P_{+}:=\begin{pmatrix}P_{0}\\ &P_{0}\\ &&\ddots\\ &&&P_{0}\end{pmatrix}\ \mbox{with }P_{0}:=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\ \mbox{and }P_{-}:=\frac{1}{N+1}P_{+}.$
Proof.
The proof follows integrating by parts. ∎
Definition 5.2.
Consider the spaces $\mathcal{J}_{+}\cong\mathcal{J}_{-}\cong\mathbb{C}^{2(N+1)}$ edowed with the indefinite non-degenerated products
$\langle x,y\rangle_{+}:=\left(P_{+}x\mid y\right)\ \mbox{and }\langle x,y\rangle_{-}=\left(P_{-}x\mid y\right).$
$(\mathcal{J}_{\pm},\langle\cdot,\cdot\rangle_{\pm})$ are Krein spaces.
Definition 5.3.
We define the function
$\begin{split}\mathbb{F}:&\ G(\mathcal{H}_{0}^{*})\to\mathcal{J}_{+}\oplus\mathcal{J}_{-}\\ &(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})\mapsto\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix},\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)\end{split}$ (5.4)
and the sesquilinear form
$\begin{split}\rho:&\ (\mathcal{J}_{+}\oplus\mathcal{J}_{-})\times(\mathcal{J}_{+}\oplus\mathcal{J}_{-})\to\mathbb{C}\\ &((x,y),(\tilde{x},\tilde{y}))\mapsto\langle x,\tilde{x}\rangle_{+}-\langle y,\tilde{y}\rangle_{-}.\end{split}$ (5.5)
From (5.2), (5.3) and the previous definition we conclude for $u\in D(\mathcal{H}_{0}^{*})$ that
$\Lambda((\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}),(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}))=\rho(\mathbb{F}(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}),\mathbb{F}(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})).$ (5.6)
It follows directly from (5.6) and [44] the following theorem.
Theorem 5.2.
Let $\mathcal{H}$ be an extension of $\mathcal{H}_{0}$. $\mathcal{H}$ is self-adjoint if and only if there exists a subspace $X\subset\mathcal{J}_{+}\oplus\mathcal{J}_{-}$, $X$ $\rho$-self-orthogonal with $\mathbb{F}(G(\mathcal{H}))=X$.
5.3. Extensions generating unitary dynamics
In a similar fashion as done to obtain Theorem 4.3 in the case of the Airy operator, we can characterize the $\rho$-self-orthogonal subspaces in terms of the graph of a linear operator that, according to Theorem 5.2, characterize the self-adjoint extensions of $\mathcal{H}_{0}$.
Theorem 5.3.
A subspace $X\subset\mathcal{J}_{+}\oplus\mathcal{J}_{-}$ is $\rho$-self-orthogonal if and only if there is $\mathcal{L}:\mathcal{J}_{+}\to\mathcal{J}_{-}$ unitary such that $X=G(\mathcal{L})$.
Proof.
The proof follows in a similar fashion as in the proof of Theorem 4.3. ∎
With our first example we show that the extensions inducing the so-called $\delta$ interactions can be seen as a member of a $\frac{(N+2)(N+1)}{2}$-parameters family of extensions generating unitary dynamics.
Example 7 ($\delta$ type interactions on looping edge graphs).
Let us first set $N=1$ and let $\mathcal{L}_{1}$ be defined by
$\mathcal{L}_{1}=\begin{pmatrix}1&0&0&0\\ m_{1}&2&m_{2}&0\\ 0&0&1&0\\ m_{2}&0&m_{3}&2\end{pmatrix}.$ (5.7)
It is not hard to see that $\mathcal{L}_{1}$ is a $(\mathcal{J}_{+},\mathcal{J}_{-})$-unitary operator. From Theorems 5.2 and 5.3 $\mathcal{L}_{1}$ defines a self-adjoint extension $\mathcal{H}_{\mathcal{L}_{1}}$ of $\mathcal{H}_{0}$ via the relation $\mathbb{F}(G(\mathcal{H}_{\mathcal{L}_{1}}))=G(\mathcal{L}_{1})$. That is, an extension $\mathcal{H}_{\mathcal{L}_{1}}$ with domain made of functions $(\phi,\psi)\in H^{2}(\mathcal{G})$ such that
$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\ \phi^{\prime}(L)-2\phi^{\prime}(-L)=(m_{1}+m_{2})\phi(-L),\ \mbox{and }\\ &\phi^{\prime}(L)-2\psi^{\prime}(L)=(m_{2}+m_{3})\phi(-L).\end{split}$ (5.8)
Adding the equations in (5.8) we get
$\phi^{\prime}(L)-\phi^{\prime}(-L)=\psi^{\prime}(L)+\frac{m_{1}+2m_{2}+m_{3}}{2}\phi(-L).$ (5.9)
Of course by taking $m_{1}=-m_{3}$ we recover $D_{Z,1}$ defined in (1.5).
More generally, let us consider $N\in\mathbb{Z}^{+}$ arbitrary and parameters $m_{ij}\in\mathbb{R}$ such that $m_{ij}=m_{ji}$ for $i,j\in\{1,\dots,N+1\}$. We define $\mathcal{L}_{N}$ as a $2(N+1)\times 2(N+1)$ matrix given in terms of $2\times 2$ blocks $(A_{ij})_{ij}$ obeying the rule
$A_{ii}=\begin{pmatrix}1&0\\ m_{ii}&N+1\end{pmatrix}\mbox{ and }A_{ij}=\begin{pmatrix}0&0\\ m_{ij}&0\end{pmatrix}\mbox{ if }i\neq j.$ (5.10)
We abuse notation to denote with $\psi_{0}(L):=\phi(-L)$ and $\psi^{\prime}_{0}(L)=\phi^{\prime}(-L)$. Note that if $i=j$ we have that
$A_{ii}\partial\psi_{i-1}(L)=\begin{pmatrix}\psi_{i-1}(L)\\ m_{ii}\psi_{i-1}(L)+(N+1)\psi^{\prime}_{i-1}(L)\end{pmatrix}$ (5.11)
and if $i\neq j$ we have
$A_{ij}\partial\psi_{j-1}(L)=\begin{pmatrix}0\\ m_{ij}\psi_{j-1}(L)\end{pmatrix}.$ (5.12)
From the structure of the rows of $\mathcal{L}_{N}$ it can be seen that it is a full-rank invertible matrix. In particular we conclude $\mathcal{L}_{N}$ is injecitve. Moreover, a tedious but elementary computation shows that $\mathcal{L}_{N}^{*}P_{-}\mathcal{L}_{N}=P_{+}$, which implies that for all $x,y\in\mathbb{C}^{2(N+1)}$
$\langle\mathcal{L}_{N}x,\mathcal{L}_{N}y\rangle_{-}=\langle x,y\rangle_{+}.$
The latter implies $\mathcal{L}_{N}^{\#}=\mathcal{L}_{N}^{-1}$, showing $\mathcal{L}_{N}$ defines a $(\mathcal{J}_{+},\mathcal{J}_{-})$-unitary operator. From Theorems 5.2 and 5.3, $\mathcal{L}_{N}$ defines a self-adjoint extension $\mathcal{H_{\mathcal{L}_{N}}}$ of $\mathcal{H}_{0}$ with domain consisting of the functions in $D(\mathcal{H}_{0}^{*})$ such that
$\mathcal{L}_{N}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{1}(L)\\ \vdots\\ \partial\psi_{N}(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(L)\\ \partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}.$ (5.13)
Note from the even rows of $\mathcal{L}_{N}$ that we are a priori fixing $\phi(-L)=\phi(L)=\psi_{1}(L)=\cdots=\psi_{N}(L)$. Taking this into account, let us open the expression (5.13). We have
$\begin{split}&A_{11}\partial\psi_{0}(L)+A_{12}\partial\psi_{1}(L)+\cdots+A_{1(N+1)}\partial\psi_{N}(L)=\partial\phi(L),\ \\ &A_{21}\partial\psi_{0}(L)+A_{22}\partial\psi_{1}(L)+\cdots+A_{2(N+1)}\partial\psi_{N}(L)=\partial\phi(L),\ \\ &\hskip 142.26378pt\vdots\\ &A_{(N+1)1}\partial\psi_{0}(L)+A_{(N+1)2}\partial\psi_{1}(L)+\cdots+A_{(N+1)(N+1)}\partial\psi_{N}(L)=\partial\phi(L),\ \\ \end{split}$ (5.14)
Adding the equations in (5.14) we get
$\sum_{i=1}^{N+1}A_{i1}\partial\psi_{0}(L)+A_{i2}\partial\psi_{1}(L)+\cdots+A_{i(N+1)}\partial\psi_{N}(L)=(N+1)\partial\phi(L).$ (5.15)
According to (5.11) and (5.12), we can rewrite (5.15) as
$\frac{1}{N+1}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N+1}\begin{pmatrix}0\\ m_{ij}\psi_{j-1}(L)\end{pmatrix}+\frac{1}{N+1}\sum_{i=1}^{N+1}\begin{pmatrix}\psi_{i-1}(L)\\ m_{ii}\psi_{i-1}(L)+(N+1)\psi_{i-1}^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi(L)\\ \phi^{\prime}(L)\end{pmatrix}.$ (5.16)
From the continuity at the vertex and the second entries in (5.16) we conclude
$\phi^{\prime}(L)-\phi^{\prime}(-L)=\sum_{i=1}^{N}\psi^{\prime}_{i}(L)+\phi(-L)\frac{1}{N+1}\sum_{i,j=1}^{N+1}m_{ij}.$ (5.17)
Any choice of the parameters such that $\frac{1}{N+1}\sum_{i,j=1}^{N+1}m_{ij}=Z\in\mathbb{R}$ makes $D(\mathcal{H}_{\mathcal{L}_{N}})=D_{Z,N}$, where $D_{Z,N}$ is defined on (1.5).
Remark 1.
It is worth mentioning that the operator $\mathcal{L}_{N}$ was built by fixing continuity at the vertex (see the even rows of $\mathcal{L}_{N}$) and removing the problematic entries that won’t allow $\mathcal{L}_{N}$ to be a real unitary matrix in its boundary system. Therefore, we can infer from the previous example that all the self-adjoint extensions of the Schrödinger operator coming from a real matrix that include continuity at the vertex and are posed on a looping-edge graph are of the form $(\mathcal{H},D(\mathcal{H}))=(\mathcal{H}_{0},D_{Z,N})$ with $Z$ varying on $\mathbb{R}$.
The following example pretend to explore families of operators that induce unitary dynamics and are compatible with continuity of the derivatives at the vertex.
Example 8 ($\delta^{\prime}$ type interactions on looping edge and $\mathcal{T}$-graphs).
Consider $N=1$ (tadpole case). At first, we begin considering $-L$ as an independent vertex, i.e., we won’t fix the condition $\phi(-L)=\phi(L)$ directly from the operator. Let $m_{1},\dots,m_{4}\in\mathbb{R}$ be parameters. Define
$\mathcal{L}=\begin{pmatrix}m_{1}&m_{2}&0&0\\ 0&\frac{2}{m_{1}}&0&m_{3}\\ -m_{1}m_{3}&-m_{2}m_{3}&2&m_{4}\\ 0&0&0&1\end{pmatrix}.$ (5.18)
It can be seen that $\mathcal{L}$ is a $(\mathcal{J}_{+},\mathcal{J}_{-})$-unitary operator. From theorems 5.2 and 5.3 we have a self-adjoint extension $\mathcal{H}_{\mathcal{L}}$ of $\mathcal{H}_{0}$ with domain made of the functions in $D(\mathcal{H}_{0}^{*})$ such that
$\begin{split}&\psi^{\prime}(L)=\phi^{\prime}(L),\ \phi(L)-m_{1}\phi(-L)=m_{2}\phi^{\prime}(-L),\ \phi^{\prime}(L)-2m_{1}^{-1}\phi^{\prime}(-L)=m_{3}\psi^{\prime}(L)\\ &\mbox{ and }\phi(L)+m_{1}m_{3}\phi(-L)=-m_{2}m_{3}\phi^{\prime}(-L)+2\psi(L)+m_{4}\psi^{\prime}(L).\end{split}$ (5.19)
In particular if we take $m_{1}=1$ and $m_{2}=0$ we recover the looping structure $\phi(-L)=\phi(L)$ since (5.19) can be rewritten as
$\begin{split}&\psi^{\prime}(L)=\phi^{\prime}(L),\ \phi(L)=\phi(-L),\ \phi^{\prime}(L)-2\phi^{\prime}(-L)=m_{3}\psi^{\prime}(L)\\ &\mbox{ and }\phi(L)+m_{3}\phi(-L)-2\psi(L)=m_{4}\psi^{\prime}(L).\end{split}$ (5.20)
Moreover, if we additionally let $m_{3}=-1$ in (5.20) we get
$\begin{split}&\psi^{\prime}(L)=\phi^{\prime}(L)=\phi^{\prime}(-L),\ \phi(L)=\phi(-L),\\ &\mbox{ and }\psi(L)=\frac{-m_{4}}{2}\psi^{\prime}(L).\end{split}$ (5.21)
Remark 2.
The existence and stability of standing waves profiles for the NLS model (1.1) on a tadpole graph, ${\bf{U}}(x,t)=e^{-i\omega t}(\phi,\psi)$, with $(\phi,\psi)$ satisfying the $\delta^{\prime}$-type interactions in (5.21) is the aim of a work in progress by the authors.
5.4. Extensions generating unitary dynamics parameterized by linear subspaces
We now emulate the strategy of separating derivatives used for the Airy operator. We note in this case there is no actual separation of derivatives but another reference frame to build the extensions. We begin by following the same approach as done in Section 4.3 to rewrite (5.3) in a simpler way.
Proposition 5.4.
For any $\textbf{U},\textbf{V}\in D(\mathcal{H}_{0}^{*})$ we have
$[\mathcal{H}_{0}^{*}\textbf{U},\textbf{V}]-[\textbf{U},\mathcal{H}_{0}^{*}\textbf{V}]=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)-\left(Q\vec{\textbf{U}}\mid\vec{\textbf{V}}^{\prime}\right)=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)-\left(\vec{\textbf{U}}\mid Q\vec{\textbf{V}}^{\prime}\right),$ (5.22)
where
$Q=\begin{pmatrix}1\\ &-1\\ &&1\\ &&&\ddots\\ &&&&1\end{pmatrix}.$
Definition 5.4.
Given a subspace $Y\subset\mathbb{C}^{2+N}$ consider the operator $\mathcal{H}_{Y}$, acting as $\mathcal{H}_{Y}\textbf{U}=\mathcal{H}_{0}\textbf{U}$, with domain
$D(\mathcal{H}_{Y}):=\left\{\textbf{U}\in D(\mathcal{H}_{0}^{*})\mid\ \vec{\textbf{U}}\in Y\ \mbox{and }Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}\right\}.$ (5.23)
The next proposition determine explicitly the adjoint operator of $\mathcal{H}_{Y}$.
Proposition 5.5.
For any closed subspace $Y\subset\mathbb{C}^{2+N}$ the operator $\mathcal{H}_{Y}$ is a self-adjoint extension of $\mathcal{H}_{0}$.
Proof.
It is enough to prove that the adjoint of $\mathcal{H}_{Y}$ is given by $\mathcal{H}_{Y}^{*}\textbf{V}=\mathcal{H}_{0}^{*}\textbf{V}$ with domain
$D(\mathcal{H}_{Y}^{*}):=\left\{\textbf{V}\in D(\mathcal{H}_{0}^{*})\mid\ \vec{\textbf{V}}\in Y\ \mbox{and }Q\vec{\textbf{V}}^{\prime}\in Y^{\perp}\right\}=D(\mathcal{H}_{Y}).$ (5.24)
It is obvious from the definition of $D(\mathcal{H}_{Y})$ that if $Q\vec{\textbf{V}}^{\prime}\in Y^{\perp}$ and $\vec{\textbf{V}}\in Y$ then $\textbf{V}\in D(\mathcal{H}_{Y}^{*})$.
Let us prove the converse. Let $\textbf{V}\in D(\mathcal{H}_{0}^{*})$. By definition V is in $D(\mathcal{H}_{Y}^{*})$ if and only if $[\mathcal{H}_{Y}\textbf{U},\textbf{V}]=[\textbf{U},\mathcal{H}_{Y}^{*}\textbf{V}]$ for all $\textbf{U}\in D(\mathcal{H}_{Y})$.
Let $x\in Y^{\perp}$ be arbitrary. There exists $\textbf{U}\in D(\mathcal{H}_{Y})$ such that $\vec{\textbf{U}}=0$ and $Q\vec{\textbf{U}}^{\prime}=x$. From (5.22), we must have
$\left(x\mid\vec{\textbf{V}}\right)=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)=0.$
Since $x$ was arbitrary, it must be the case $\vec{\textbf{V}}\in(Y^{\perp})^{\perp}=Y$.
Similarly, for $x\in Y$ there exists $\textbf{U}\in D(\mathcal{H}_{Y})$ such that $\vec{\textbf{U}}^{\prime}=0$ and $\vec{\textbf{U}}=x$. From (5.22) we conclude it must be the case
$\left(x\mid Q\vec{\textbf{V}}^{\prime}\right)=\left(\vec{\textbf{U}}\mid Q\vec{\textbf{V}}^{\prime}\right)=0.$
We therefore have $Q\vec{\textbf{V}}^{\prime}\in Y^{\perp}$. ∎
Let us give some examples on the tadpole graph. Assume $N=1$.
Example 9.
Consider the space $Y_{0}=\mbox{span}\{(1,1,0)^{T}\}$. Note $\vec{\textbf{U}}\in Y_{0}$ if and only if $\phi(-L)=\phi(L)$. Also, $Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}$ if and only if $\phi^{\prime}(-L)=\phi^{\prime}(L)$. We therefore have that
$D(\mathcal{H}_{Y_{0}})=\{u\in D(\mathcal{H}_{0}^{*})\mid\ \phi(-L)=\phi(L)\ \mbox{and }\phi^{\prime}(L)=\phi^{\prime}(-L)\}.$
Example 10.
Consider the space $Y=\mbox{span}\{(1,0,-1)\}$. Consequently, $Y^{\perp}=\{(x,y,w)\mid x=w\}$. Note $\vec{\textbf{U}}\in Y$ if and only if $\phi(L)=0$ and $\phi(-L)=-\psi(L)$. Similarly, $Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}$ if and only if $\phi^{\prime}(-L)=\psi^{\prime}(L)$. According to Proposition 5.5, ${\mathcal{H}}_{Y}$ defines a self-adjoint extension of $\mathcal{H}_{0}$ with domain
$D(\mathcal{H}_{Y})=\{\textbf{U}\in D(\mathcal{H}_{0}^{*})\mid\phi(-L)=-\psi(L),\ \phi(L)=0,\ \mbox{and}\ \phi^{\prime}(-L)=\psi^{\prime}(L)\}.$ (5.25)
5.5. The Schrödinger operator on $\mathcal{T}$-shaped metric graphs
As highlighted in the introduction of the section, the analysis performed in the previous subsections for the Schrödinger equation on looping-edge graphs extends to $\mathcal{T}$-shaped graphs, i.e., to graphs $\mathcal{G}$ that consists on an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with 00 (see Figure 4). A similar frame-work can be employed to obtain self-adjoint extensions of $\mathcal{H}_{0}$ posed on these kind of metric graphs. We proceed to exhibit some examples of them.
$L$
00
Figure 4. $\mathcal{T}$-shaped metric graph.
Example 11.
Let $N=1$ (exactly one half-line attached to 00) and consider again the operator $\mathcal{L}$ defined in (5.18) with $m_{1}=1$ and $m_{2}=0$ (see (5.20)). Instead of $m_{3}=-1$, let us now consider $m_{3}=1$. It can be obtained analogously to (5.20) an extension of $\mathcal{H}_{0}$ with domain of functions in $H^{2}(\mathcal{G})$ such that
$\begin{split}&\psi^{\prime}(0+)=\phi^{\prime}(0-),\ \phi^{\prime}(-L)=0,\ \phi(0-)=\phi(-L)\ \\ &\mbox{ and }\phi(0-)-\psi(0+)=\frac{m_{4}}{2}\psi^{\prime}(0+),\end{split}$ (5.26)
where $0+$ and $0-$ refer to the right and left-handed limits respectively.
Thus, (5.26) is an example of the kind of domain of extensions that might be related to interactions of $\delta^{\prime}$-type in a $\mathcal{T}$-graph.
With the following example we intend to present, under this frame of reference, all the extensions generating unitary dynamics with continuity in the derivatives at the vertex. This family includes the so-called $\delta^{\prime}$ interactions. For simplicity we work with $N=1$ and $N=2$.
Example 12.
Let us first set $N=1$ and let $\mathcal{L}$ be defined by
$\mathcal{L}=\begin{pmatrix}2&m_{1}&0&m_{2}\\ 0&1&0&0\\ 0&m_{2}&2&m_{3}\\ 0&0&0&1\end{pmatrix}.$ (5.27)
It can be seen that in a respective boundary system $\mathcal{L}$ is a unitary operator which defines a self-adjoint extension $\mathcal{H}_{\mathcal{L}}$ of $\mathcal{H}_{0}$ via the relation $\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(0+)\end{pmatrix}=\begin{pmatrix}\partial\phi(0-)\\ \partial\phi(0-)\end{pmatrix}$. That is, an extension $\mathcal{H}_{\mathcal{L}}$ with domain made of functions $(\phi,\psi)\in H^{2}(\mathcal{G})$ such that
$\begin{split}&\phi^{\prime}(-L)=\phi^{\prime}(0-)=\psi^{\prime}(0+),\ \phi(0+)-2\phi(-L)=(m_{1}+m_{2})\psi^{\prime}(0+)\mbox{ and }\\ &\phi(0-)-2\psi(0+)=(m_{2}+m_{3})\psi^{\prime}(0+).\end{split}$ (5.28)
Adding the last two equations in (5.28) we get
$\phi(0-)-\phi(-L)=\psi(0+)+\frac{m_{1}+2m_{2}+m_{3}}{2}\psi^{\prime}(0+).$ (5.29)
In particular, if $m_{1}=-m_{3}$ it follows
$\phi(0-)-\phi(-L)=\psi(0+)+m_{2}\psi^{\prime}(0+).$ (5.30)
Similarly, let us assume $N=2$ and consider the operator
$\mathcal{L}_{2}=\begin{pmatrix}3&m_{1}&0&m_{2}&0&m_{3}\\ 0&1&0&0&0&0\\ 0&m_{2}&3&m_{4}&0&m_{5}\\ 0&0&0&1&0&0\\ 0&m_{3}&0&m_{5}&3&m_{6}\\ 0&0&0&0&0&1\end{pmatrix}$ (5.31)
It can be seen that $\mathcal{L}_{2}$ defines a unitary extension $H_{\mathcal{L}_{2}}$ of $\mathcal{H}_{0}$ with domain consisting of the functions $(\phi,\psi_{1},\psi_{2})\in H^{2}(\mathcal{G})$ such that
$\begin{split}&\phi^{\prime}(-L)=\phi^{\prime}(0-)=\psi^{\prime}_{1}(0+)=\psi^{\prime}_{2}(0+),\ \phi(0-)-3\phi(-L)=(m_{1}+m_{2}+m_{3})\psi^{\prime}(0+)\\ &\phi(0-)-3\psi_{1}(0+)=(m_{2}+m_{4}+m_{5})\psi^{\prime}(0+)\mbox{ and }\phi(0-)-3\psi_{2}(0+)=(m_{3}+m_{5}+m_{6})\psi^{\prime}(0+).\end{split}$ (5.32)
Adding the last three equations we get
$\phi(0-)-\phi(-L)=\sum_{j=1}^{2}\psi_{j}(0+)+(m_{1}+m_{4}+m_{6}+2m_{2}+2m_{3}+2m_{5})\psi^{\prime}(0+).$ (5.33)
Of course there is a choice of the parameters that turn (5.33) into
$\phi(0-)-\phi(-L)=\sum_{j=1}^{2}\psi_{j}(0+)+Z\psi^{\prime}(0+),\ \ Z\in\mathbb{R}.$ (5.34)
Moreover, the operator $\mathcal{L}_{2}$ can be generalized to any amount of half-lines $N\in\mathbb{Z}^{+}$ by following a similar procedure to the one done in Example 7. That way it can be obtained an extension with domain made of functions satisfying continuity in the derivatives and
$\phi(0-)-\phi(-L)=\sum_{j=1}^{N}\psi_{j}(0+)+Z\psi^{\prime}(0+),\ \ Z\in\mathbb{R}.$ (5.35)
Acknowledgments
A. Muñoz is partially supported by CNPq, Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brazil grant 170246/2023-0 and INCTMat, Instituto Nacional de Ciência e Tecnologia de Matemática - Brazil. J. Angulo was partially funded by CNPq/Brazil Grant.
Data availability statement. The data that support the findings of this study are available upon request from the authors.